### v1

# AlphaNet: Scaling Up Local Frame-based Atomistic Foundation Model

Abstract
We present AlphaNet, a local frame-based equivariant model designed to achieve both accurate and efficient simulations for atomistic systems. Recently, machine learning force fields (MLFFs) have gained prominence in molecular dynamics simulations due to their advantageous efficiency-accuracy balance compared to classical force fields and quantum mechanical calculations, alongside their transferability across various systems. Despite the advancements in improving model accuracy, the efficiency and scalability of MLFFs remain significant obstacles in practical applications. AlphaNet enhances computational efficiency and accuracy by leveraging the local geometric structures of atomic environments through the construction of equivariant local frames and learnable frame transitions. We substantiate the efficacy of AlphaNet across diverse datasets, including defected graphene, formate decomposition, zeolites, and surface reactions. AlphaNet consistently surpasses well-established models, such as NequIP and DeepPot, in terms of both energy and force prediction accuracy. Notably, AlphaNet offers one of the best trade-offs between computational efficiency and accuracy among existing models. Moreover, AlphaNet exhibits scalability across a broad spectrum of system and dataset sizes, affirming its versatility. Our code and data is available at https://github.com/zmyybc/AlphaNet.
00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors. Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com, haixiao@tsinghua.edu.cn and henkelman@utexas.edu.
1 Introduction
Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems [24, 3, 22]. While atomic forces necessary for these simulations can theoretically be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limit their application to small systems and short time scales. As a result, many phenomena occurring over longer time frames and larger spatial scales remain inaccessible, even with the most powerful supercomputers.
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended periods [7, 23, 17]. However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics.
The advent of machine learning has introduced a promising solution to this dilemma. By training models on data derived from ab initio calculations, machine learning forcefields (MLFF) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields. Unlike traditional models that rely on explicit functional forms to describe bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
As interest in applying machine learning techniques to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 [35], is a large-scale pre-trained model with improved attention architecture over DeepPot [32]. Similarly, JMP [28] leverages diverse molecular systems of different types as a joint pre-training model. MACE [1], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of machine learning-driven atomistic simulations [13, 2, 26, 29, 8].
However, the balance between computational efficiency and accuracy is key to the practical application of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computation more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks which accuracy is more important.
Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetry including rotation, translation and reflection. To reduce the unnecessary degree of freedoms, equivariant models are developed to build model that respects these symmetries  [25, 31]. Among all the developed models, most of the expressive models by date [2, 1, 21] are based on spherical harmonics [15]. However, the computation to calculate the tensor product of irreducible representations impose expensive computational overheads. While another branch of work which achieve rotation equivariance through building equivariant frames which could either be local or global [18, 9, 33, 12]. The main benefit of frame-based approaches is that is eliminates the necessity to use tensor product which greatly improves the computational efficiency.
In this paper, we propose a local frame-based atomistic equivariant model (AlphaNet) for accurate yet highly efficient atomistic system simulations. Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance frame transition and temporal connection for multi-scale modeling. Extensive quantitative experiments demonstrate AlphaNet excels at accuracy, efficiency and scalability compared to state-of-the-art MLFF models on a variety of molecular systems, from defected graphene, formate decomposition, zeolites, to surface reactions.
2 Experiment
2.1 Dataset
The Defected Bilayer Graphene Dataset consists of reference structures specifically designed to train and validate machine learning potentials (MLPs). It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers). The dataset comprises single-point DFT (PBE+MBD) energies and atomic forces calculated for various configurations, including different interlayer distances, stacking modes, and manually deformed structures. Additionally, snapshot configurations from MD simulations at different temperatures were included. The data were split into training, validation, and test sets, containing 3988, 4467, and 200 structures, respectively. Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.  [34]
The Formate Decomposition on Cu Dataset consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond. The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase. Nudged elastic band (NEB) method [16] was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations. These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory. AlphaNet was trained on 2500 uniformly sampled structures from the full dataset, with a validation set of 250 structures and the mean absolute error evaluated on the remaining structures.  [2]
The Zeolite Dataset comprises essential reference structures designed for various applications such as catalysis, adsorption, and separation. Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability. The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) [4] simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) [19, 20]. For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration. The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
The OC2M Dataset is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials [5]. This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both). Each of them contains around 1 million structures. It encompasses configurations involving 56 different chemical elements, ensuring a broad and diverse representation of atomic environments. These configurations were generated using density functional theory (DFT) calculations, with a focus on a wide range of materials science applications, including bulk materials, surfaces, and defect structures. Originating from the comprehensive OC20 Dataset, the OC2M subset provides accurate total energies and atomic forces for each configuration. This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
Figure 1: Illustration of Experimental Systems: a) The Formate Decomposition on Cu Dataset. Cu atoms are depicted in brown, C atoms in grey, O atoms in red, and H atoms in white. b) The Zeolite Dataset. Li atoms are shown in purple, Si atoms in grey, Al atoms in orange, and O atoms in red. c) The Defected Bilayer Graphene Dataset.
2.2 Results
2.2.1 Accuracy
Our model demonstrates excellent precision across various experimental datasets. On datasets such as formate decomposition and defected graphene, as detailed in Table 1. We selected Nequip [2] as our baseline model due to its demonstrated superior performance in the research associated with the dataset. AlphaNet consistently outperforms the baseline model in terms of both energy and force accuracy.
The formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface. For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Å for force and 0.23 meV/atom for energy, compared to Nequip’s 47.3 meV/Å and 0.50 meV/atom, respectively. These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules. This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
Another challenge for machine learning force fields (MLFFs) is accurately modeling the sliding layer effect in materials like defected graphene, where interlayer interactions are critical. On the defected graphene test set, AlphaNet achieves a force MAE of 32.0 meV/Å and an energy MAE of 1.7 meV/atom, substantially outperforming Nequip, which records 60.2 meV/Å for force MAE and 1.9 meV/atom for energy MAE. Furthermore, AlphaNet successfully reproduces the binding energy curve for AB-stacked bilayer graphene as calculated by PBE+MBD, without requiring explicit long-range dispersion interaction corrections. The deviation in the shallow sliding potential energy landscape is less than 0.4 meV/atom, illustrating AlphaNet’s capacity to handle systems with subtle interlayer forces and complex structural dynamics (as shown in Figure 3).
This ability to model both metallic and covalent bonding scenarios, as well as interlayer interactions, makes AlphaNet a versatile and powerful tool for materials design, particularly in fields requiring high accuracy for large-scale systems. These results suggest that AlphaNet can be effectively applied to a wide variety of systems, ranging from catalytic surfaces to layered materials, where precision in both energy and force calculations is critical for predicting material behavior.
2.2.2 Scaling
In the context of machine learning force field (MLFF) models, performance across large and diverse datasets is crucial for validating model robustness and scalability. The zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations, provides an excellent benchmark to evaluate the accuracy of these models. Here, we compare the performance of Deep Potential [32] and AlphaNet using two key metrics: Energy MAE (meV/atom) and Force MAE (meV/Å). A lower mean absolute error (MAE) indicates higher predictive accuracy.
Table 4 shows the results of Deep Pot and AlphaNet across different zeolite configurations. Across nearly all tested configurations, AlphaNet significantly outperforms Deep Pot in both energy and force predictions.
For instance, in the ABWopt configuration, AlphaNet achieves an Energy MAE of 41 meV, substantially lower than Deep Pot’s 90 meV. Similarly, AlphaNet records a Force MAE of 54 meV/Å, outperforming Deep Pot’s 90 meV/Å. This pattern of superiority is consistent across most configurations. Notable results include the CANopt structure, where AlphaNet shows an exceptional Force MAE of 10 meV/Å compared to Deep Pot’s 90 meV/Å, indicating a marked improvement in force prediction accuracy. Additionally, in the FERopt structure, AlphaNet achieves an Energy MAE of 100 meV, significantly lower than Deep Pot’s 290 meV, and a Force MAE of 18 meV/Å, much better than Deep Pot’s 130 meV/Å.
The performance of the model in terms of Mean Absolute Error (MAE) for both energy (in meV) and force (in meV/Å) improves as the training data size increases from 5k to 80k samples. Specifically, as shown in Figure 2, the energy MAE decreases from 220 meV at 5k samples to 41 meV at 80k samples. Similarly, the force MAE decreases from 241 meV/Å at 5k samples to 54 meV/Å at 80k samples. This trend highlights the scalable performance of increasing the training dataset size in enhancing the accuracy of both energy and force predictions on the zeolite dataset.
(a) Energy MAE (meV) vs Data Size
(b) Force MAE (meV/Å) vs Data Size
Figure 2: MAE of Energy and Force predictions as a function of training data size on the zeolite dataset.
Even in cases where Deep Pot performs competitively, such as in the LOSopt configuration, where Deep Pot achieves an Energy MAE of 110 meV/atom compared to AlphaNet’s 130 meV, AlphaNet still maintains an advantage in Force MAE with a value of 46 meV/Å compared to Deep Pot’s 70 meV/Å. This suggests that while Deep Pot may occasionally perform well in energy prediction, AlphaNet is more consistent in providing better overall results, particularly in force predictions.
In terms of standout performance, AlphaNet demonstrates its predictive power in configurations like LTJopt, where it achieves an Energy MAE of just 1.2 meV compared to Deep Pot’s 70 meV, and in the GISopt structure, where it delivers a Force MAE of only 3.1 meV/Å, far surpassing Deep Pot’s 50 meV/Å.
These results indicate that AlphaNet is better equipped to handle complex zeolite configurations, providing both lower energy and force prediction errors across most datasets. This demonstrates its robustness and makes it a promising model for large-scale applications in materials design, where accuracy and scalability are critical.
The Open Catalyst Project (OCP) provides the OC20 dataset, a comprehensive collection of materials and structures designed to train machine learning models for catalysis research. At present, our work has focused on scaling our model to the OC2M S2EF task, with plans to extend the training to the full OC20 dataset in future research. We compared the performance of our model against established models such as EquiformerV2 [21], SchNet [26], and DimeNet++ [14].
We were able to compare our results with SchNet and DimeNet++ models trained on the full OC20 dataset ( 134M data points), as well as with EquiformerV2 trained on both the full dataset and the smaller 2M subset. Currently we report our results on the validation set as it is public available. Results on the test set would be done as soon as possible. It’s important to note that each of these models was trained under different configurations, including training for both energy and force, training for energy only, and training for force only. The corresponding performance metrics are summarized in Table 2.
To control the training cost and make sure all of the elements are seen in the data, we first train our model on the full OC training set for 10,000 steps and then turns to 2M subset. We believe this would not provide additional benefits because we have tested that this procedure makes no difference in the validation ID set as there is no additional elements here. Our model trained on 2M subset for 2,000,000 steps demonstrates exceptional accuracy in energy prediction, achieving a mean absolute error (MAE) of 0.19 eV, which is significantly better than EquiformerV2’s performance on both the full dataset (0.229 eV) and the 2M subset (0.285 eV). SchNet and DimeNet++, which were trained on the full dataset, exhibit MAE values above 0.35 eV. This trend holds consistently across different test sets, demonstrating the robustness of our model in energy prediction.
In terms of force prediction, while our model trained on the 2M subset does not surpass larger models like EquiformerV2 and DimeNet++-large, it still achieves competitive results relative to models of a similar scale trained on the full dataset. Table 3 provides additional insights, revealing that our model has substantially lower training costs and fewer parameters compared to EquiformerV2 and DimeNet++. This efficiency, combined with the model’s high precision on this large dataset, indicates that our approach is computationally effective, requiring less data and resources while still delivering strong performance.
Given these promising results, we are confident that our model can be successfully scaled to the full OC20 dataset in future work, further enhancing its predictive accuracy for both energy and force calculations. This would allow us to continue improving the state-of-the-art in machine learning for catalysis research.
2.2.3 Speed
The primary advantage of machine learning force fields (MLFFs) over density functional theory (DFT) force fields lies in their significantly faster inference speed. While increasing the number of parameters in an ML model can improve its precision, it often results in slower inference times. To evaluate the speed of different MLFFs, we measured the average forward pass time across 2,000 zeolite structures with a batch size of 10. We also tested systems of varying sizes by expanding the unit cells. Although we aimed to maintain a consistent number of parameters across models, slight variations exist due to differences in architecture.
We compared several graph-based models, including Nequip, MACE, SchNet, and PaiNN, against our AlphaNet. As illustrated in Figure 4, AlphaNet shows a distinct speed advantage over other MLFFs, with a relatively slow increase in computational time as system size grows. This demonstrates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
Table 1: Performance of Models on Different Datasets
Dataset Metric NequIP AlphaNet
Formate Decomposition Force MAE (meV/Å) 47.3 45.5
Energy MAE (meV/atom) 0.50 0.23
Defected Graphene Force MAE (meV/Å) 60.2 32.0
Energy MAE (meV/atom) 1.9 1.7
Table 2: Energy and Force MAE for different models
Model ID OOD-adsorbate OOD-catalyst OOD-both
Energy MAE (eV) ↓
SchNet-full 0.43 0.49 0.53 0.72
SchNet-full-energy-only 0.38 0.45 0.55 0.70
SchNet-full-force-only 34.00 33.77 35.30 38.47
DimeNet++-full 0.48 0.47 0.53 0.65
DimeNet++-full-energy-only 0.35 0.40 0.51 0.65
DimeNet++-full-force-only 28.20 28.94 28.91 34.90
EquiformerV2-2M($\lambda_{E}=2$) 0.28 - - -
EquiformerV2-full($\lambda_{E}=2$) 0.22 - - -
AlphaNet-2M($\lambda_{E}=4$) 0.19 0.23 0.23 0.29
AlphaNet-2M($\lambda_{E}=20$) 0.18 0.23 0.25 0.30
Force MAE (eV/Å) ↓
SchNet-full 0.050 0.054 0.052 0.064
SchNet-full-energy-only 0.485 0.496 0.485 0.547
SchNet-full-force-only 0.045 0.048 0.047 0.058
DimeNet++-full 0.045 0.047 0.045 0.055
DimeNet++-full-energy-only 0.341 0.339 0.341 0.362
DimeNet++-full-force-only 0.034 0.035 0.035 0.041
EquiformerV2($\lambda_{E}=2$) 0.022 - - -
EquiformerV2-full($\lambda_{E}=2$) 0.016 - - -
AlphaNet-2M($\lambda_{E}=4$) 0.036 0.037 0.035 0.040
AlphaNet-2M($\lambda_{E}=20$) 0.054 0.062 0.061 0.070
Note: The downward arrow (↓) indicates that lower values are preferable. The term ’full’ in a model’s name signifies that it was trained on the complete OC20 training set, ’2M’ indicates that the model was trained on OC2M sub set. $\lambda_{E}$ refers to the weight of the energy loss relative to the force, with a value set to 100. Bold indicates the best performance, while underline represents the second best.
Table 3: Training Performance and Training Time for Different Models
Models Energy MAE (eV) ↓ Force MAE (eV/Å) ↓ Training Time (GPU Days) ↓ Number of Parameters ↓
SchNet 0.549 0.0568 194 9.1M
DimeNet++ 0.515 0.0328 1600 10.1M
EquiformerV2 0.236 0.0157 1368 153M
AlphaNet 0.201 0.0356 360 6.1M
Figure 3: Comparisons between AlphaNet predictions and DFT calculations. Parity plots for (a) total energies and (b) atomic forces obtained for the training (blue), validation (orange), and test (green) data sets.(c) AlphaNet (middle) and DFT (left) sliding PES for bilayer graphene, and their difference (right).
Table 4: Performance of Deep Pot and AlphaNet on the Zeolite Dataset
Dataset Metric Deep Pot AlphaNet
ABWopt Energy MAE 90 71
Force MAE 90 54
BCTopt Energy MAE 110 73
Force MAE 50 55
BPHopt Energy MAE 210 41
Force MAE 60 90
CANopt Energy MAE 150 62
Force MAE 90 10
EDIopt Energy MAE 40 5.4
Force MAE 50 59
FERopt Energy MAE 290 100
Force MAE 130 18
GISopt Energy MAE 60 56
Force MAE 50 3.1
JBWopt Energy MAE 150 18
Force MAE 70 50
LOSopt Energy MAE 110 130
Force MAE 70 46
LTAopt Energy MAE 150 90
Force MAE 64 60
LTJopt Energy MAE 70 1.2
Force MAE 110 60
NATopt Energy MAE 210 47
Force MAE 110 9.9
PARopt Energy MAE 90 6.5
Force MAE 70 3.5
PHIopt Energy MAE 60 16
Force MAE 120 50
SODopt Energy MAE 200 9.5
Force MAE 110 21
THOopt Energy MAE 160 70
Force MAE 60 47
Figure 4: Comparison of inference speed for various machine learning force fields. The x-axis represents the number of atoms in the systems, while the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches.
3 Discussion
Alphanet presents a significant advancement as a scalable machine learning potential, offering both high precision and fast computational speed. This success can be attributed to its local-frame-based architecture, which scalarizes geometric quantities over local frames and effectively aggregate local frames to construct local- and global-aware features. By utilizing this architecture, our model excels at capturing complex interactions across a broad spectrum of systems, including metallic, covalent, and ionic bonding, as well as long-range interactions.
One of the key advantages of Alphanet lies in its scalability. Specifically, our model is capable of maintaining high performance as the size of the dataset increases, which sets it apart from other large-scale atomistic models. Despite the growing complexity introduced by larger datasets, Alphanet operates with fewer parameters, making it more efficient while still delivering robust results. This ability to handle large datasets without sacrificing accuracy positions Alphanet as a promising solution for extensive simulations and real-world applications. In future work, we aim to extend the model’s capabilities to explore additional types of interactions, such as hydrogen bonding, which are critical in a wide range of chemical and biological systems.
Another avenue for further exploration is the interpretability of the model. While our current implementation achieves impressive results, gaining a deeper understanding of how the model interprets and processes interactions at the atomic level will enhance its applicability and trustworthiness in practical scenarios.
Despite these advancements, one challenge remains: the efficient use of GPU memory during training, particularly when dealing with large systems consisting of over 1,000 atoms (though we can parallelize it across multiple GPUs). This is a common limitation shared by most message-passing neural networks (MPNNs), and addressing it will be crucial for enabling the model to scale further in both size and complexity. Future work will focus on optimizing memory usage to overcome this bottleneck and ensure the model’s applicability to even larger systems.
4 Methodology
4.1 Problem Formulation
Machine learning forcefields. A single state of a molecular system can be described by a set of atom types $h\in\mathbb{R}^{n\times d}$ and atomic positions $x\in\mathbb{R}^{n\times 3}$. In molecular dynamics, we are interested in learning a function $f:\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}$ which predicts the energy of the current state of the molecular system readily to be used for simulation. The force can be further calculated by $-\nabla_{x}f(x)$, thus often referred as to machine learning force fields (MLFFs).
Message-passing neural networks. One common way to model molecular systems is through message-passing neural networks, which is a general neural architecture that leverages both permutation equivariance and locality in graph-structured data. Specifically, there are two essential operations in message-passing neural networks, message calculation and message update. The message calculation operation encodes structured messages from local environments $m_{i}=\oplus_{j\in\mathcal{N}(i)}f_{m}(f_{h}(x_{i}),h(x_{j}),e_{ij})$ where $i$ and $j$ denotes the index of different nodes, $e_{ij}$ denotes optional edge features and $f_{m}$ and $f_{h}$ denotes neural networks that encode both edge and node features; The message update aggregates the incoming messages from the neighbors $m_{i}=f_{u}(m_{i})$, where $f_{u}$ is a neural network.
Equivariant message-passing neural networks. Equivariance is another crucial property to building MLFFs as molecular systems are invariant to the Euclidean group. A function $f$ is said to be equivariant with respect to certain group $G$ if $f\circ g(x)=g\circ f(x)$, $\forall x\in X,\,g\in G$. One natural property of message-passing neural networks is the permutation equivariance such that changing the order of the input set of nodes do not result in difference output for the same node as it only depends on the neighbors which are unchanged under permutation. In addition to permutation, we need to consider the special Euclidean group in 3D, SE(3), which includes rotation and translation operations. We do so by leveraging local frames [10]. Specifically, we build a set of equivariant and complete frames and scalarize the geometric quantity which does not break SE(3)-equivariance under nonlinear neural network encodings.
To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following [10]
$\mathcal{F}_{ij}=\left(\hat{e}^{1}_{ij},\hat{e}^{2}_{ij},\hat{e}^{3}_{ij}\right)=\left(\frac{x_{i}-x_{j}}{\|x_{i}-x_{j}\|},\ \frac{(x_{i}-\overline{x})\times(x_{j}-\overline{x})}{\|(x_{i}-\overline{x})\times(x_{j}-\overline{x})\|},\ \frac{(x_{i}-x_{j})\times\left((x_{i}-\overline{x})\times(x_{j}-\overline{x})\right)}{\|(x_{i}-x_{j})\times\left((x_{i}-\overline{x})\times(x_{j}-\overline{x})\right)\|}\right)$
To scalarize geometric quantity such as a vector $v$, we take the inner product between it and the frames which results in a set of invariant scalars and can be transformed back through a tensorization operation
$\text{Scalarize}(v,\mathcal{F}_{ij})=(s^{1}_{ij},s^{2}_{ij},s^{3}_{ij})=(e^{1}_{ij}\cdot v,e^{2}_{ij}\cdot v,e^{3}_{ij}\cdot v)$
$\text{Tensorize}(s_{ij},\mathcal{F}_{ij})=s^{1}_{ij}e^{1}_{ij}+s^{2}_{ij}e^{2}_{ij}+s^{3}_{ij}e^{3}_{ij}$
Efficient and expressive equivariant message-passing neural networks. Following [11], we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames. We consider two modules to improve the expressiveness while maintaining the efficiency of the network. Specifically, we leverage a local structure encoding module which scalarizes not only features from local neighbors but also from overlapping neighbors along each edge, denoted as $A_{ij}:=f_{l}(\text{Scalarize}(S_{i-j},\mathcal{F}_{ij}))$, where $f_{l}$ denotes a neural network and $S_{i-j}$ represents overlapping neighbors of node $i$ and $j$. In addition, we use an additional frame transition module which encodes the alignment between each neighboring frame to improve the expressiveness. We realize this by incorporating vector-based message passing to amortize the cost where $\mathbb{m}_{ij}$ denotes equivariant edge features. This framework is highly efficient as it avoids expensive higher-body message-passing neural networks and higher-order tensor updates.
$\{h_{i},x_{i}\}=\{f_{u}(\oplus_{j\in\mathcal{N}(i)}f_{m}(h(x_{i}),h(x_{j}),A_{ij},e_{ij})),\oplus_{j\in\mathcal{N}(i)}f_{\mathbb{m}}(h(x_{i}),h(x_{j}),A_{ij},e_{ij})\mathbb{m}_{ij}\}$
Rotary positional embedding as invariant frame transition. Positional embedding are known to improve and stabilize the training of transformers [27]. We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture [30]. In fact the frame transition can be considered as a generalized rotary positional embedding for 3D equivariant frames. Note that the spirit of rotary position embedding is to design an embedding scheme for relative position in sequential data. To build the rotary positional embedding,  [30] utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
$\langle f_{q}(x_{i},i),f_{k}(x_{j},j)\rangle=g(x_{i},x_{j},i-j)$
where $f_{q}$ and $f_{k}$ denote the neural network embeddings of query and key inside an attention layer. Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices [30]:
$\begin{pmatrix}\cos\theta_{n},&-\sin\theta_{n}\\ \sin\theta_{n},&\cos\theta_{n}\end{pmatrix}$
where $\theta_{n}=10000^{-2\frac{(n-1)}{d}},\quad n\in[1,2,\ldots,\frac{d}{2}]$, and $d$ is the dimension of features. From a geometric point of view, the index $i$ of $x_{i}$ provides a one-dimensional position embedding such that the inner product $\langle f_{q}(x_{i},i),f_{k}(x_{j},j)\rangle$ only depend on the relative position $i-j$. In the three-dimensional world, we can utilize the frame transition matrices between $x_{i}$ and $x_{j}$ that plays a similar rule of the index difference $i-j$ in a one-dimensional sequence. In addition, we introduce the rotary positional embedding on invariant features $h$ and represent the rotation by the multiplication of complex numbers. We predict an additional complex number from our invariant features $h_{i}$ for each node $i$ and multiply it by the scalarized features $h_{i}$. Similar to the equivariant message passing in each layer of the network, the learned rotary embedding is also applied in each layer.
Temporal connection. In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network [6]. As message-passing neural networks are local, we believe it is beneficial to integrate multi-scale information. Specifically, for each consecutive layer, we learn a kernel that linear transforms and aggregates features. Given the invariant feature embeddings $h^{l_{1}}$ at layer $l_{1}$ and $h^{(l_{2})}$ at layer $l_{2}$, the kernel is of the matrix product form: $M^{(l_{3,1,2})}$ which transforms $h^{(l_{1})}$ and $h^{(l_{2})}$ to a new feature $h^{(l_{3})}$ and added as a residual into the next layer:
$h^{(l_{3})}=\sum_{k_{1}=1}^{K_{1}}\biggl{(}\sum_{k_{2}=1}^{K_{2}}M^{(l_{3,1,2})}_{k_{1},k_{2}}\cdot h^{(l_{2})}_{k_{2}}\biggl{)}\cdot\,h^{(l_{1})}_{k_{1}}$
[…]

### v2

@@
-AlphaNet: Scaling Up Local Frame-based Atomistic Foundation Model
+AlphaNet: Scaling Up Local Frame-based Neural Interatomic Potential
 Abstract
@@
 00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors.
-Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com, haixiao@tsinghua.edu.cn and henkelman@utexas.edu.
+Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com and haixiao@tsinghua.edu.cn.
 1 Introduction
 Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems [30, 4, 24].
-While atomic forces necessary for these simulations can theoretically be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limit their application to small systems and short time scales.
-As a result, many phenomena occurring over longer time frames and larger spatial scales remain inaccessible, even with the most powerful supercomputers.
-Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended periods [7, 23, 17].
+While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps).
+As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
+Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories [8, 25, 19].
 However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics.
-The advent of machine learning has introduced a promising solution to this dilemma.
-By training models on data derived from ab initio calculations, machine learning forcefields (MLFF) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields.
-Unlike traditional models that rely on explicit functional forms to describe bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
-As interest in applying machine learning techniques to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models.
+The advent of machine learning techniques has introduced a promising solution to this dilemma.
+By training models on data derived from first-principles calculations, machine learning force fields (MLFFs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields.
+Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
+As interest in applying MLFFs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models.
 One example among the others, DPA-1 [43], is a large-scale pre-trained model with improved attention architecture over DeepPot [40].
@@
 MACE [1], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions.
-Collectively, these models represent significant strides in the evolution of machine learning-driven atomistic simulations [13, 2, 26, 29, 8].
-However, the balance between computational efficiency and accuracy is key to the practical application of these models.
-On one side, efficient yet less expressive models are more suitable for tasks that weigh computation more than accuracy.
-On the other side, less efficient yet expressive models are more suitable for tasks which accuracy is more important.
-Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetry including rotation, translation and reflection.
-To reduce the unnecessary degree of freedoms, equivariant models are developed to build model that respects these symmetries  [25, 31].
-Among all the developed models, most of the expressive models by date [2, 1, 21] are based on spherical harmonics [15].
-However, the computation to calculate the tensor product of irreducible representations impose expensive computational overheads.
-While another branch of work which achieve rotation equivariance through building equivariant frames which could either be local or global [18, 9, 33, 12].
-The main benefit of frame-based approaches is that is eliminates the necessity to use tensor product which greatly improves the computational efficiency.
-In this paper, we propose a local frame-based atomistic equivariant model (AlphaNet) for accurate yet highly efficient atomistic system simulations.
-Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance frame transition and temporal connection for multi-scale modeling.
-Extensive quantitative experiments demonstrate AlphaNet excels at accuracy, efficiency and scalability compared to state-of-the-art MLFF models on a variety of molecular systems, from defected graphene, formate decomposition, zeolites, to surface reactions.
+Collectively, these models represent significant strides in the evolution of MLFF-driven atomistic simulations [14, 2, 33, 36, 9].
+Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models.
+On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy.
+On the other side, less efficient yet expressive models are more suitable for tasks where accuracy is more important.
+Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection.
+Thus, equivariant models are developed to respect these symmetries  [32, 38], and most of these expressive models by date [2, 1, 23] are based on the spherical harmonics [17], but calculating the tensor product of irreducible representations imposes expensive computational overheads.
+While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global [20, 10, 41, 13], and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
+In this paper, we propose a local-frame-based atomistic equivariant model, named AlphaNet, for accurate yet highly efficient atomistic simulations.
+Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling.
+Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art MLFF models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, to surface reactions.
 2 Experiment
@@
 This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
+The MPTraj Dataset and Matbench Discovery Test.
+MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials [9].
+Matbench Discovery [31] is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set [39].
+In this case we train our model on the MPTraj dataset and test it on the WBM test set.
 Figure 1: Illustration of Experimental Systems: a) The Formate Decomposition on Cu Dataset.
@@
 The Open Catalyst Project (OCP) provides the OC20 dataset, a comprehensive collection of materials and structures designed to train machine learning models for catalysis research.
-At present, our work has focused on scaling our model to the OC2M S2EF task, with plans to extend the training to the full OC20 dataset in future research.
-We compared the performance of our model against established models such as EquiformerV2 [21], SchNet [26], and DimeNet++ [14].
-We were able to compare our results with SchNet and DimeNet++ models trained on the full OC20 dataset ( 134M data points), as well as with EquiformerV2 trained on both the full dataset and the smaller 2M subset.
-Currently we report our results on the validation set as it is public available.
-Results on the test set would be done as soon as possible.
-It’s important to note that each of these models was trained under different configurations, including training for both energy and force, training for energy only, and training for force only.
+In our current work, we focus on scaling our model to the OC2M S2EF task, with plans to extend training to the full OC20 dataset in future research.
+We compare the performance of our model against established benchmarks, including EquiformerV2 [23], EScAIP [29], eSCN [28], GemNet-OC [16], SchNet [33], and DimeNet++ [15].
+Our comparisons include SchNet and DimeNet++ models trained on the full OC20 dataset ($\sim$134M data points), as well as EquiformerV2, EScAIP, eSCN, and GemNet-OC trained on the smaller 2M subset.
+Specifically, we utilize pretrained models of SchNet and DimeNet++ available at https://fair-chem.github.io/core/model_checkpoints.html, while results for the other models are sourced from [29].
+All reported results are based on the validation set, as it is publicly accessible.
 The corresponding performance metrics are summarized in Table 2.
-To control the training cost and make sure all of the elements are seen in the data, we first train our model on the full OC training set for 10,000 steps and then turns to 2M subset.
-We believe this would not provide additional benefits because we have tested that this procedure makes no difference in the validation ID set as there is no additional elements here.
-Our model trained on 2M subset for 2,000,000 steps demonstrates exceptional accuracy in energy prediction, achieving a mean absolute error (MAE) of 0.19 eV, which is significantly better than EquiformerV2’s performance on both the full dataset (0.229 eV) and the 2M subset (0.285 eV).
-SchNet and DimeNet++, which were trained on the full dataset, exhibit MAE values above 0.35 eV.
-This trend holds consistently across different test sets, demonstrating the robustness of our model in energy prediction.
-In terms of force prediction, while our model trained on the 2M subset does not surpass larger models like EquiformerV2 and DimeNet++-large, it still achieves competitive results relative to models of a similar scale trained on the full dataset.
-Table 3 provides additional insights, revealing that our model has substantially lower training costs and fewer parameters compared to EquiformerV2 and DimeNet++.
-This efficiency, combined with the model’s high precision on this large dataset, indicates that our approach is computationally effective, requiring less data and resources while still delivering strong performance.
+To manage training costs and ensure exposure to all elements in the dataset, we first train our model on the full OC training set for 10,000 steps before switching to the 2M subset.
+We note that this initial step does not provide additional benefits, as testing confirms no significant difference in performance on the validation ID set, which contains no additional elements.
+Our model, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV.
+This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit higher MAE values above 0.35 eV.
+This trend remains consistent across various test sets, highlighting the robustness of our model in energy prediction.
+In terms of force prediction, while our model trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP, it remains competitive relative to models of similar scale trained on the full dataset.
+This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$.
+Table 3 provides an efficiency evaluation of different models.
+Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than models like EScAIP.
+It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for autograd, resulting in higher time and memory costs.
+Speak to WBM test, AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4.
+With an F1 score of 0.789 and DAF score of 4.315, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints.
+Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$).
+This combination of efficiency and physics-aware design positions AlphaNet as a robust foundation for scalable materials modeling.
 Given these promising results, we are confident that our model can be successfully scaled to the full OC20 dataset in future work, further enhancing its predictive accuracy for both energy and force calculations.
@@
 The primary advantage of machine learning force fields (MLFFs) over density functional theory (DFT) force fields lies in their significantly faster inference speed.
-While increasing the number of parameters in an ML model can improve its precision, it often results in slower inference times.
+While increasing the number of parameters in an ML model can enhance its precision, this often comes at the cost of slower inference times.
 To evaluate the speed of different MLFFs, we measured the average forward pass time across 2,000 zeolite structures with a batch size of 10.
-We also tested systems of varying sizes by expanding the unit cells.
+Additionally, we tested systems of varying sizes by expanding the unit cells.
 Although we aimed to maintain a consistent number of parameters across models, slight variations exist due to differences in architecture.
 We compared several graph-based models, including NequIP, MACE, SchNet, and PaiNN, against our AlphaNet.
-As illustrated in Figure 4, AlphaNet shows a distinct speed advantage over other MLFFs, with a relatively slow increase in computational time as system size grows.
-This demonstrates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
+As illustrated in Figure 4 (a), AlphaNet demonstrates a distinct speed advantage over other MLFFs, exhibiting a relatively slow increase in computational time as system size grows.
+This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
+Furthermore, we evaluated the inference speed of pretrained models on the Matbench Discovery leaderboard.
+To ensure fairness, we selected models that are compliant, open-source, and accessible without additional requests.
+We measured the average speed of energy prediction using the same API.
+The models included AlphaNet, ORB-v2 [26], SevenNet [27], GRACE [3], MACE, and CHGNet.
+Figure 4 (b) presents the results.
+Among these models, GRACE, which is based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based.
+Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-0 and SevenNet-l3i5).
+This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.
 Table 1: Performance of Models on Different Datasets
@@
 Energy MAE (meV/atom) 1.9 1.7
-Table 2: Energy and Force MAE for different models
-Model ID OOD-adsorbate OOD-catalyst OOD-both
-Energy MAE (eV) ↓
-SchNet-full 0.43 0.49 0.53 0.72
-SchNet-full-energy-only 0.38 0.45 0.55 0.70
-SchNet-full-force-only 34.00 33.77 35.30 38.47
-DimeNet++-full 0.48 0.47 0.53 0.65
-DimeNet++-full-energy-only 0.35 0.40 0.51 0.65
-DimeNet++-full-force-only 28.20 28.94 28.91 34.90
-EquiformerV2-2M($\lambda_{E}=2$) 0.28 - - -
-EquiformerV2-full($\lambda_{E}=2$) 0.22 - - -
-AlphaNet-2M($\lambda_{E}=4$) 0.19 0.23 0.23 0.29
-AlphaNet-2M($\lambda_{E}=20$) 0.18 0.23 0.25 0.30
-Force MAE (eV/Å) ↓
-SchNet-full 0.050 0.054 0.052 0.064
-SchNet-full-energy-only 0.485 0.496 0.485 0.547
-SchNet-full-force-only 0.045 0.048 0.047 0.058
-DimeNet++-full 0.045 0.047 0.045 0.055
-DimeNet++-full-energy-only 0.341 0.339 0.341 0.362
-DimeNet++-full-force-only 0.034 0.035 0.035 0.041
-EquiformerV2($\lambda_{E}=2$) 0.022 - - -
-EquiformerV2-full($\lambda_{E}=2$) 0.016 - - -
-AlphaNet-2M($\lambda_{E}=4$) 0.036 0.037 0.035 0.040
-AlphaNet-2M($\lambda_{E}=20$) 0.054 0.062 0.061 0.070
+Table 2: Results on the OC20 validation set
+Model Parameters ↓ Energy MAE (eV) ↓ Force MAE (eV/Å) ↓
+SchNet-full 9.1M 0.54 0.55
+DimeNet++-full 10.1M 0.53 0.048
+GemNet-OC 38M 0.29 0.026
+eSCN 51M 0.28 0.021
+EScAIP-Medium 146M 0.25 0.019
+EquiformerV2($\lambda_{E}=2$) 146M 0.28 0.022
+AlphaNet($\lambda_{E}=4$) 6.1M 0.25 0.040
+AlphaNet($\lambda_{E}=20$) 6.1M 0.24 0.062
 Note: The downward arrow (↓) indicates that lower values are preferable.
-The term ’full’ in a model’s name signifies that it was trained on the complete OC20 training set, ’2M’ indicates that the model was trained on OC2M sub set.
+The term ’full’ in a model’s name signifies that it was trained on the complete OC20 training set, the others are trained on the 2M subset.
 $\lambda_{E}$ refers to the weight of the energy loss relative to the force, with a value set to 100.
-Bold indicates the best performance, while underline represents the second best.
-Table 3: Training Performance and Training Time for Different Models
-Models Energy MAE (eV) ↓ Force MAE (eV/Å) ↓ Training Time (GPU Days) ↓ Number of Parameters ↓
-SchNet 0.549 0.0568 194 9.1M
-DimeNet++ 0.515 0.0328 1600 10.1M
-EquiformerV2 0.236 0.0157 1368 153M
-AlphaNet 0.201 0.0356 360 6.1M
+Bold indicates the best performance.
+Table 3: Efficiency on OC dataset, measured on A100 40GB GPU
+Models Parameters Training Speed (Sample/Sec) ↑ Training Memory (GB/Sample) ↓ Inference Speed (Sample/Sec) ↑ Inference Memory (GB/Sample) ↓ Conserv.
+SchNet 9.1M 39 2.5 112.1 1.7
+DimeNet++ 10.1M 29 3.6 92 2.5
+GemNet-OC 38M 16.22 1.63 39.2 1.61 $\times$
+eSCN 51M 8.5 1.75 25.2 1.74 $\times$
+EScAIP 83M 37.8 1.23 111.8 1.09 $\times$
+EquiformerV2 31M 15.7 1.63 36.72 1.61 $\times$
+AlphaNet 6.1M 41.2 2.2 120.1 1.5 $\checkmark$
+Table 4: Results on WBM full test set
+Model F1 ↑ DAF ↑ Prec ↑ Acc ↑ MAE ↓ R2 ↑ Size Type
+eqV2 S DeNS 0.798 4.362 0.748 0.927 0.035 0.785 31.2M EFSD
+AlphaNet (ours) 0.789 4.315 0.740 0.923 0.039 0.742 16.2M EFSG
+DPA3-v1-MPtrj 0.757 4.13 0.709 0.911 0.04 0.793 3.37M EFSG
+ORB MPtrj 0.755 4.188 0.719 0.911 0.043 0.752 25.2M EFSD
+SevenNet-l3i5 0.751 4.112 0.706 0.909 0.042 0.773 1.17M EFSG
+SevenNet-0 0.719 3.804 0.653 0.893 0.046 0.75 842k EFSG
+GRACE-2L-MPtrj 0.687 3.714 0.637 0.884 0.05 0.74 15.3M EFSG
+MACE-MP-0 0.668 3.4 0.583 0.867 0.055 0.698 4.69M EFSG
+CHGNet 0.612 3.038 0.521 0.839 0.061 0.69 413k EFSGM
+ [^1] Data source:Open Catalyst https://fair-chem.github.io/core/model_checkpoints.html; Matbench Discovery https://matbench-discovery.materialsproject.org/
 Figure 3: Comparisons between AlphaNet predictions and DFT calculations.
@@
 Figure 4: Comparison of inference speed for various machine learning force fields.
-The x-axis represents the number of atoms in the systems, while the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches.
+(a) Inference speed as a function of system size.
+The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches.
+(b) Inference speed and F1 performance of models accessible and compliant with the Matbench Discovery leaderboard.
+The speed is calculated as the number of energies predicted per second.
 3 Discussion
@@
 Future work will focus on optimizing memory usage to overcome this bottleneck and ensure the model’s applicability to even larger systems.
+Improving the smoothness of the output is also crucial, as we believe it would enhance the models’ generalization ability and yield a more physically meaningful potential energy surface.
 4 Methodology

### v3

@@
-AlphaNet: Scaling Up Local Frame-based Neural Interatomic Potential
+AlphaNet: Scaling Up Local Frame-based Neural Network Interatomic Potentials
 Abstract
-We present AlphaNet, a local frame-based equivariant model designed to achieve both accurate and efficient simulations for atomistic systems.
-Recently, machine learning force fields (MLFFs) have gained prominence in molecular dynamics simulations due to their advantageous efficiency-accuracy balance compared to classical force fields and quantum mechanical calculations, alongside their transferability across various systems.
-Despite the advancements in improving model accuracy, the efficiency and scalability of MLFFs remain significant obstacles in practical applications.
-AlphaNet enhances computational efficiency and accuracy by leveraging the local geometric structures of atomic environments through the construction of equivariant local frames and learnable frame transitions.
-We substantiate the efficacy of AlphaNet across diverse datasets, including defected graphene, formate decomposition, zeolites, and surface reactions.
-AlphaNet consistently surpasses well-established models, such as NequIP and DeepPot, in terms of both energy and force prediction accuracy.
-Notably, AlphaNet offers one of the best trade-offs between computational efficiency and accuracy among existing models.
-Moreover, AlphaNet exhibits scalability across a broad spectrum of system and dataset sizes, affirming its versatility.
+Molecular dynamics simulations demand unprecedented accuracy and scalability to tackle grand challenges in energy materials, catalytic processes, and biomolecular design.
+To bridge this gap, we present AlphaNet, a local frame-based equivariant model that simultaneously advances computational efficiency and predictive precision for atomistic systems.
+By constructing equivariant local frames with learnable geometric transitions, AlphaNet encodes atomic environments with enhanced representational capacity, achieving state-of-the-art accuracy in energy and force predictions.
+Extensive benchmarks—spanning defected graphene, formate decomposition, inorganic bulks, and large-scale datasets (OC2M and Matbench Discovery) demonstrate its superior performance over existing neural network interatomic potentials while ensuring scalability across diverse system sizes.
+The synergy of accuracy, efficiency, and transferability positions AlphaNet as a transformative tool for simulating multiscale phenomena, from catalyst dynamics to energy storage interfaces, with direct implications for accelerating the discovery of functional materials and complex molecular systems.
 Our code and data is available at https://github.com/zmyybc/AlphaNet.
 00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors.
-Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com and haixiao@tsinghua.edu.cn.
+Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@gmail.com and haixiao@tsinghua.edu.cn.
 1 Introduction
-Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems [30, 4, 24].
+Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems (Richards et al., 2016; Boero et al., 1998; Lindorff-Larsen et al., 2011).
 While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps).
 As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
-Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories [8, 25, 19].
-However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics.
+Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories (Jones and Chapman, 1924; Cornell et al., 1995; MacKerell et al., 1998).
+However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics (Iftimie et al., 2005).
 The advent of machine learning techniques has introduced a promising solution to this dilemma.
-By training models on data derived from first-principles calculations, machine learning force fields (MLFFs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields.
-Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
-As interest in applying MLFFs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models.
-One example among the others, DPA-1 [43], is a large-scale pre-trained model with improved attention architecture over DeepPot [40].
-Similarly, JMP [35] leverages diverse molecular systems of different types as a joint pre-training model.
-MACE [1], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions.
-Collectively, these models represent significant strides in the evolution of MLFF-driven atomistic simulations [14, 2, 33, 36, 9].
+By training models on data derived from first-principles calculations, neural network interatomic potentials (NNIPs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields.
+Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, NNIPs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
+As interest in applying NNIPs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models.
+One example among the others, DPA-1 (Zhang et al., 2024a), is a large-scale pre-trained model with improved attention architecture over DeepPot (Wang et al., 2018).
+Similarly, JMP (Shoghi et al., 2024) leverages diverse molecular systems of different types as a joint pre-training model.
+MACE (Batatia et al., 2022), on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions.
+Collectively, these models represent significant strides in the evolution of NNIP-driven atomistic simulations (Qu and Krishnapriyan, 2024; Liao et al., 2024; Gasteiger et al., 2021; Batzner et al., 2022; Schütt et al., 2018; Smith et al., 2017; Deng et al., 2023).
 Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models.
@@
 Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection.
-Thus, equivariant models are developed to respect these symmetries  [32, 38], and most of these expressive models by date [2, 1, 23] are based on the spherical harmonics [17], but calculating the tensor product of irreducible representations imposes expensive computational overheads.
-While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global [20, 10, 41, 13], and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
-In this paper, we propose a local-frame-based atomistic equivariant model, named AlphaNet, for accurate yet highly efficient atomistic simulations.
-Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling.
-Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art MLFF models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, to surface reactions.
-2 Experiment
-2.1 Dataset
-The Defected Bilayer Graphene Dataset consists of reference structures specifically designed to train and validate machine learning potentials (MLPs).
-It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers).
-The dataset comprises single-point DFT (PBE+MBD) energies and atomic forces calculated for various configurations, including different interlayer distances, stacking modes, and manually deformed structures.
-Additionally, snapshot configurations from MD simulations at different temperatures were included.
-The data were split into training, validation, and test sets, containing 3988, 4467, and 200 structures, respectively.
-Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.
-[42]
-The Formate Decomposition on Cu Dataset consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond.
-The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase.
-Nudged elastic band (NEB) method [18] was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations.
-These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory.
-AlphaNet was trained on 2500 uniformly sampled structures from the full dataset, with a validation set of 250 structures and the mean absolute error evaluated on the remaining structures.
-[2]
-The Zeolite Dataset comprises essential reference structures designed for various applications such as catalysis, adsorption, and separation.
-Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability.
-The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) [5] simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) [21, 22].
-For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration.
-The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
-The OC2M Dataset is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials [6].
-This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both).
-Each of them contains around 1 million structures.
-It encompasses configurations involving 56 different chemical elements, ensuring a broad and diverse representation of atomic environments.
-These configurations were generated using density functional theory (DFT) calculations, with a focus on a wide range of materials science applications, including bulk materials, surfaces, and defect structures.
-Originating from the comprehensive OC20 Dataset, the OC2M subset provides accurate total energies and atomic forces for each configuration.
-This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
-The MPTraj Dataset and Matbench Discovery Test.
-MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials [9].
-Matbench Discovery [31] is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set [39].
-In this case we train our model on the MPTraj dataset and test it on the WBM test set.
-Figure 1: Illustration of Experimental Systems: a) The Formate Decomposition on Cu Dataset.
-Cu atoms are depicted in brown, C atoms in grey, O atoms in red, and H atoms in white.
-b) The Zeolite Dataset.
-Li atoms are shown in purple, Si atoms in grey, Al atoms in orange, and O atoms in red.
-c) The Defected Bilayer Graphene Dataset.
-2.2 Results
-2.2.1 Accuracy
-Our model demonstrates excellent precision across various experimental datasets.
-On datasets such as formate decomposition and defected graphene, as detailed in Table 1.
-We selected NequIP [2] as our baseline model due to its demonstrated superior performance in the research associated with the dataset.
-AlphaNet consistently outperforms the baseline model in terms of both energy and force accuracy.
-The formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface.
-For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Å for force and 0.23 meV/atom for energy, compared to NequIP’s 47.3 meV/Å and 0.50 meV/atom, respectively.
+Thus, equivariant models are developed to respect these symmetries  (Satorras et al., 2021; Thomas et al., 2018), and most of these expressive models by date (Batzner et al., 2022; Batatia et al., 2022; Liao et al., 2024) are based on the spherical harmonics (Geiger and Smidt, 2022), but calculating the tensor product of irreducible representations imposes expensive computational overheads.
+While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global (Jumper et al., 2021; Du et al., 2022, 2023; Duval et al., 2023), and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
+In this paper, we propose a local-frame-based equivariant atomistic model, named AlphaNet, for accurate yet highly efficient atomistic simulations.
+Building on the success of frame-based NNIPs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling.
+Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art (SOTA) NNIP models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, surface reactions and inorganic bulks.
+2 Results
+We systematically assess the capabilities of AlphaNet by examining its prediction accuracy, scalability with respect to model and dataset size, and computational efficiency in terms of inference speed and memory usage.
+Initially, we validate atomic-level prediction accuracy across five diverse chemical systems, simultaneously demonstrating AlphaNet’s scalability concerning dataset and system sizes.
+Subsequently, we quantify computational efficiency by benchmarking both the memory footprint and inference costs of the model.
+Comprehensive details regarding datasets, comparative methodologies, and hyperparameter configurations are provided in Appendix B.
+2.1 AlphaNet achieving high accuracy across all datasets.
+We first consider the formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface.
+For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Åfor force and 0.23 meV/atom for energy, compared to NequIP (Batzner et al., 2022) 47.3 meV/Åand 0.50 meV/atom, respectively, as shown in Table 1.
 These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules.
 This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
-Another challenge for machine learning force fields (MLFFs) is accurately modeling the sliding layer effect in materials like defected graphene, where interlayer interactions are critical.
-On the defected graphene test set, AlphaNet achieves a force MAE of 32.0 meV/Å and an energy MAE of 1.7 meV/atom, substantially outperforming NequIP, which records 60.2 meV/Å for force MAE and 1.9 meV/atom for energy MAE.
+Another challenge for NNIP models is accurately modeling the sliding layer effect in materials like defected graphene, where inter-layer interactions are critical.
+On the defected graphene dataset, AlphaNet achieves a force MAE of 32.0 meV/Åand an energy MAE of 1.7 meV/atom, substantially outperforming NequIP, which records 60.2 meV/Åfor force MAE and 1.9 meV/atom for energy MAE.
 Furthermore, AlphaNet successfully reproduces the binding energy curve for AB-stacked bilayer graphene as calculated by PBE+MBD, without requiring explicit long-range dispersion interaction corrections.
 The deviation in the shallow sliding potential energy landscape is less than 0.4 meV/atom, illustrating AlphaNet’s capacity to handle systems with subtle interlayer forces and complex structural dynamics (as shown in Figure 1).
-This ability to model both metallic and covalent bonding scenarios, as well as interlayer interactions, makes AlphaNet a versatile and powerful tool for materials design, particularly in fields requiring high accuracy for large-scale systems.
-These results suggest that AlphaNet can be effectively applied to a wide variety of systems, ranging from catalytic surfaces to layered materials, where precision in both energy and force calculations is critical for predicting material behavior.
-2.2.2 Scaling
-In the context of machine learning force field (MLFF) models, performance across large and diverse datasets is crucial for validating model robustness and scalability.
-The zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations, provides an excellent benchmark to evaluate the accuracy of these models.
-Here, we compare the performance of Deep Potential [40] and AlphaNet using two key metrics: Energy MAE (meV/atom) and Force MAE (meV/Å).
-A lower mean absolute error (MAE) indicates higher predictive accuracy.
-Table 5 shows the results of Deep Pot and AlphaNet across different zeolite configurations.
+Table 1: Results on the Formate Decomposition and Defected Graphene datasets. Boldface indicates the best performance.
+Dataset Metric NequIP AlphaNet
+Formate Decomposition Force MAE (meV/Å) $\downarrow$ 47.3 45.5
+Energy MAE (meV/atom) $\downarrow$ 0.50 0.23
+Defected Graphene Force MAE (meV/Å) $\downarrow$ 60.2 32.0
+Energy MAE (meV/atom) $\downarrow$ 1.9 1.7
+(a) Parity plot for total energies obtained for the training (blue), validation (orange), and test (green) data sets.
+(b) Parity plot for atomic forces obtained for the training (blue), validation (orange), and test (green) data sets.
+(c) AlphaNet (middle) and DFT (left) sliding PES for bilayer graphene, and their difference (right).
+Figure 1: Comparisons between AlphaNet predictions and DFT calculations on the Zeolite dataset.
+The third dataset we consider is the zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations.
+Table 2 shows the results of Deep Pot (Wang et al., 2018) and AlphaNet across different zeolite configurations.
 Across nearly all tested configurations, AlphaNet significantly outperforms Deep Pot in both energy and force predictions.
@@
 Similarly, AlphaNet records a Force MAE of 54 meV/Å, outperforming Deep Pot’s 90 meV/Å.
-This pattern of superiority is consistent across most configurations.
-Notable results include the CANopt structure, where AlphaNet shows an exceptional Force MAE of 10 meV/Å compared to Deep Pot’s 90 meV/Å, indicating a marked improvement in force prediction accuracy.
-Additionally, in the FERopt structure, AlphaNet achieves an Energy MAE of 100 meV, significantly lower than Deep Pot’s 290 meV, and a Force MAE of 18 meV/Å, much better than Deep Pot’s 130 meV/Å.
-The performance of the model in terms of Mean Absolute Error (MAE) for both energy (in meV) and force (in meV/Å) improves as the training data size increases from 5k to 80k samples.
-Specifically, as shown in Figure 2, the energy MAE decreases from 220 meV at 5k samples to 41 meV at 80k samples.
-Similarly, the force MAE decreases from 241 meV/Å at 5k samples to 54 meV/Å at 80k samples.
-This trend highlights the scalable performance of increasing the training dataset size in enhancing the accuracy of both energy and force predictions on the zeolite dataset.
-(a) Energy MAE (meV) vs Data Size
-(b) Force MAE (meV/Å) vs Data Size
-Figure 2: MAE of Energy and Force predictions as a function of training data size on the zeolite dataset.
-Even in cases where Deep Pot performs competitively, such as in the LOSopt configuration, where Deep Pot achieves an Energy MAE of 110 meV/atom compared to AlphaNet’s 130 meV, AlphaNet still maintains an advantage in Force MAE with a value of 46 meV/Å compared to Deep Pot’s 70 meV/Å.
-This suggests that while Deep Pot may occasionally perform well in energy prediction, AlphaNet is more consistent in providing better overall results, particularly in force predictions.
-In terms of standout performance, AlphaNet demonstrates its predictive power in configurations like LTJopt, where it achieves an Energy MAE of just 1.2 meV compared to Deep Pot’s 70 meV, and in the GISopt structure, where it delivers a Force MAE of only 3.1 meV/Å, far surpassing Deep Pot’s 50 meV/Å.
-These results indicate that AlphaNet is better equipped to handle complex zeolite configurations, providing both lower energy and force prediction errors across most datasets.
-This demonstrates its robustness and makes it a promising model for large-scale applications in materials design, where accuracy and scalability are critical.
-The Open Catalyst Project (OCP) provides the OC20 dataset, a comprehensive collection of materials and structures designed to train machine learning models for catalysis research.
-In our current work, we focus on scaling our model to the OC2M S2EF task, with plans to extend training to the full OC20 dataset in future research.
-We compare the performance of our model against established benchmarks, including EquiformerV2 [23], EScAIP [29], eSCN [28], GemNet-OC [16], SchNet [33], and DimeNet++ [15].
-Our comparisons include SchNet and DimeNet++ models trained on the full OC20 dataset ($\sim$134M data points), as well as EquiformerV2, EScAIP, eSCN, and GemNet-OC trained on the smaller 2M subset.
-Specifically, we utilize pretrained models of SchNet and DimeNet++ available at https://fair-chem.github.io/core/model_checkpoints.html, while results for the other models are sourced from [29].
-All reported results are based on the validation set, as it is publicly accessible.
-The corresponding performance metrics are summarized in Table 2.
-To manage training costs and ensure exposure to all elements in the dataset, we first train our model on the full OC training set for 10,000 steps before switching to the 2M subset.
-We note that this initial step does not provide additional benefits, as testing confirms no significant difference in performance on the validation ID set, which contains no additional elements.
-Our model, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV.
-This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit higher MAE values above 0.35 eV.
-This trend remains consistent across various test sets, highlighting the robustness of our model in energy prediction.
-In terms of force prediction, while our model trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP, it remains competitive relative to models of similar scale trained on the full dataset.
-This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$.
-Table 3 provides an efficiency evaluation of different models.
-Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than models like EScAIP.
-It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for autograd, resulting in higher time and memory costs.
-Speak to WBM test, AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4.
-With an F1 score of 0.789 and DAF score of 4.315, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints.
-Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$).
-This combination of efficiency and physics-aware design positions AlphaNet as a robust foundation for scalable materials modeling.
-Given these promising results, we are confident that our model can be successfully scaled to the full OC20 dataset in future work, further enhancing its predictive accuracy for both energy and force calculations.
-This would allow us to continue improving the state-of-the-art in machine learning for catalysis research.
-2.2.3 Speed
-The primary advantage of machine learning force fields (MLFFs) over density functional theory (DFT) force fields lies in their significantly faster inference speed.
-While increasing the number of parameters in an ML model can enhance its precision, this often comes at the cost of slower inference times.
-To evaluate the speed of different MLFFs, we measured the average forward pass time across 2,000 zeolite structures with a batch size of 10.
-Additionally, we tested systems of varying sizes by expanding the unit cells.
-Although we aimed to maintain a consistent number of parameters across models, slight variations exist due to differences in architecture.
-We compared several graph-based models, including NequIP, MACE, SchNet, and PaiNN, against our AlphaNet.
-As illustrated in Figure 4 (a), AlphaNet demonstrates a distinct speed advantage over other MLFFs, exhibiting a relatively slow increase in computational time as system size grows.
-This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
-Furthermore, we evaluated the inference speed of pretrained models on the Matbench Discovery leaderboard.
-To ensure fairness, we selected models that are compliant, open-source, and accessible without additional requests.
-We measured the average speed of energy prediction using the same API.
-The models included AlphaNet, ORB-v2 [26], SevenNet [27], GRACE [3], MACE, and CHGNet.
-Figure 4 (b) presents the results.
-Among these models, GRACE, which is based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based.
-Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-0 and SevenNet-l3i5).
-This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.
-Table 1: Performance of Models on Different Datasets
-Dataset Metric NequIP AlphaNet
-Formate Decomposition Force MAE (meV/Å) 47.3 45.5
-Energy MAE (meV/atom) 0.50 0.23
-Defected Graphene Force MAE (meV/Å) 60.2 32.0
-Energy MAE (meV/atom) 1.9 1.7
-Table 2: Results on the OC20 validation set
-Model Parameters ↓ Energy MAE (eV) ↓ Force MAE (eV/Å) ↓
-SchNet-full 9.1M 0.54 0.55
-DimeNet++-full 10.1M 0.53 0.048
-GemNet-OC 38M 0.29 0.026
+Table 2: Results on the Zeolite dataset. Boldface indicates the best performance. (Unit for energy is meV/atom and unit for force is meV/Å).
+ABWopt BCTopt BPHopt CANopt
+Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
+Energy MAE $\downarrow$ 90 71 110 73 210 41 150 62
+Force MAE $\downarrow$ 90 54 50 55 60 90 90 10
+EDIopt FERopt GISopt JBWopt
+Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
+Energy MAE $\downarrow$ 40 5.4 290 100 60 56 150 18
+Force MAE $\downarrow$ 50 59 130 18 50 3.1 70 50
+LOSopt LTAopt LTJopt NATopt
+Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
+Energy MAE $\downarrow$ 110 130 150 90 70 1.2 210 47
+Force MAE $\downarrow$ 70 46 64 60 110 60 110 9.9
+PARopt PHIopt SODopt THOopt
+Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
+Energy MAE $\downarrow$ 90 6.5 60 16 200 9.5 160 70
+Force MAE $\downarrow$ 70 3.5 120 50 110 21 60 47
+Table 3: Results on the OC20 validation set.
+“full” denotes models trained on the full OC20 training set, others are trained on the 2M subset.
+$\lambda_{E}$ refers to the weight of the energy loss, the weight of the force loss $\lambda_{F}$ is set to $100$.
+Boldface indicates the best performance.
+Model Parameters $\downarrow$ Energy MAE (eV) $\downarrow$ Force MAE (eV/Å) $\downarrow$
+SchNet (full) 9.1M 0.54 0.55
+DimeNet++ (full) 10.1M 0.53 0.048
+GemNet 38M 0.29 0.026
 eSCN 51M 0.28 0.021
@@
 AlphaNet ($\lambda_{E}=20$) 6.1M 0.24 0.062
-Note: The downward arrow (↓) indicates that lower values are preferable.
-The term ’full’ in a model’s name signifies that it was trained on the complete OC20 training set, the others are trained on the 2M subset.
-$\lambda_{E}$ refers to the weight of the energy loss relative to the force, with a value set to 100.
-Bold indicates the best performance.
-Table 3: Efficiency on OC dataset, measured on A100 40GB GPU
-Models Parameters Training Speed (Sample/Sec) ↑ Training Memory (GB/Sample) ↓ Inference Speed (Sample/Sec) ↑ Inference Memory (GB/Sample) ↓ Conserv.
-SchNet 9.1M 39 2.5 112.1 1.7
-DimeNet++ 10.1M 29 3.6 92 2.5
-GemNet-OC 38M 16.22 1.63 39.2 1.61 $\times$
-eSCN 51M 8.5 1.75 25.2 1.74 $\times$
-EScAIP 83M 37.8 1.23 111.8 1.09 $\times$
-EquiformerV2 31M 15.7 1.63 36.72 1.61 $\times$
-AlphaNet 6.1M 41.2 2.2 120.1 1.5 $\checkmark$
-Table 4: Results on WBM full test set
-Model F1 ↑ DAF ↑ Prec ↑ Acc ↑ MAE ↓ R2 ↑ Size Type
-eqV2 S DeNS 0.798 4.362 0.748 0.927 0.035 0.785 31.2M EFSD
-AlphaNet (ours) 0.789 4.315 0.740 0.923 0.039 0.742 16.2M EFSG
-DPA3-v1-MPtrj 0.757 4.13 0.709 0.911 0.04 0.793 3.37M EFSG
-ORB MPtrj 0.755 4.188 0.719 0.911 0.043 0.752 25.2M EFSD
-SevenNet-l3i5 0.751 4.112 0.706 0.909 0.042 0.773 1.17M EFSG
-SevenNet-0 0.719 3.804 0.653 0.893 0.046 0.75 842k EFSG
-GRACE-2L-MPtrj 0.687 3.714 0.637 0.884 0.05 0.74 15.3M EFSG
-MACE-MP-0 0.668 3.4 0.583 0.867 0.055 0.698 4.69M EFSG
-CHGNet 0.612 3.038 0.521 0.839 0.061 0.69 413k EFSGM
- [^1] Data source:Open Catalyst https://fair-chem.github.io/core/model_checkpoints.html; Matbench Discovery https://matbench-discovery.materialsproject.org/
-Figure 3: Comparisons between AlphaNet predictions and DFT calculations.
-Parity plots for (a) total energies and (b) atomic forces obtained for the training (blue), validation (orange), and test (green) data sets.(c) AlphaNet (middle) and DFT (left) sliding PES for bilayer graphene, and their difference (right).
-Table 5: Performance of Deep Pot and AlphaNet on the Zeolite Dataset
-Dataset Metric Deep Pot AlphaNet
-ABWopt Energy MAE 90 71
-Force MAE 90 54
-BCTopt Energy MAE 110 73
-Force MAE 50 55
-BPHopt Energy MAE 210 41
-Force MAE 60 90
-CANopt Energy MAE 150 62
-Force MAE 90 10
-EDIopt Energy MAE 40 5.4
-Force MAE 50 59
-FERopt Energy MAE 290 100
-Force MAE 130 18
-GISopt Energy MAE 60 56
-Force MAE 50 3.1
-JBWopt Energy MAE 150 18
-Force MAE 70 50
-LOSopt Energy MAE 110 130
-Force MAE 70 46
-LTAopt Energy MAE 150 90
-Force MAE 64 60
-LTJopt Energy MAE 70 1.2
-Force MAE 110 60
-NATopt Energy MAE 210 47
-Force MAE 110 9.9
-PARopt Energy MAE 90 6.5
-Force MAE 70 3.5
-PHIopt Energy MAE 60 16
-Force MAE 120 50
-SODopt Energy MAE 200 9.5
-Force MAE 110 21
-THOopt Energy MAE 160 70
-Force MAE 60 47
-Figure 4: Comparison of inference speed for various machine learning force fields.
+The next challenge, we consider the Open Catalyst Project OC20 dataset (Chanussot et al., 2021), a comprehensive collection of crystal structures for training NNIP models in catalysis.
+We focus on the subset OC2M and the Structure-to-energy-and-force (S2EF) task.
+We compare the performance of our model against established benchmarks, including EquiformerV2 (Liao et al., 2024), EScAIP (Qu and Krishnapriyan, 2024), eSCN (Passaro and Zitnick, 2023), GemNet-OC (Gasteiger et al., 2022), SchNet (Schütt et al., 2018), and DimeNet++ (Gasteiger et al., 2020).
+AlphaNet, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV.
+This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit worse MAE values (>0.35 eV).
+However, AlphaNet trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP in the force prediction task.
+This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$.
+Table 4: Results on the Matbench Discovery WBM unique prototype dataset. Boldface indicates the best performance.
+Model F1 $\uparrow$ DAF $\uparrow$ Prec $\uparrow$ Acc $\uparrow$ MAE $\downarrow$ R2 $\uparrow$ Size $\downarrow$ Type
+CHGNet 0.613 3.361 0.514 0.851 0.063 0.689 413k EFSGM
+MACE-MP-0 0.669 3.777 0.577 0.878 0.057 0.697 4.69M EFSG
+GRACE-2L-MPtrj 0.691 4.163 0.636 0.896 0.052 0.741 15.3M EFSG
+SevenNet-0 0.724 4.252 0.65 0.904 0.048 0.75 842k EFSG
+SevenNet-l3i5 0.76 4.629 0.708 0.92 0.044 0.776 1.17M EFSG
+ORB MPtrj 0.765 4.702 0.719 0.922 0.045 0.756 25.2M EFSD
+DPA3-v1-MPtrj 0.765 4.654 0.711 0.921 0.042 0.798 3.37M EFSG
+eqV2 S DeNS 0.815 5.042 0.771 0.941 0.036 0.788 31.2M EFSD
+EScAIP-Small 0.782 5.634 0.712 0.939 0.038 0.783 45M EFSD
+AlphaNet 0.799 4.863 0.743 0.933 0.041 0.745 16.2M EFSG
+For the last challenge, we validate the effectiveness of AlphaNet on the well-established Matbench discovery WBM test set (Riebesell et al., 2024), AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4.
+With a 0.799 F1 score and 4.863 DAF score, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints.
+Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$).
+This combination of efficiency and physics-aware design positions AlphaNet as a robust NNIP model for scalable materials modeling.
+2.2 AlphaNet is scalable—effective scaling with both model, data and system sizes
+In addition to maintaining high performance with large datasets and large systems, we validate the scaling of model size and dataset size on the zeolite dataset.
+We can observe the performance improves as the training data size increases from 50k to 800k samples, as shown in Figure 2.
+Specifically, the energy MAE decreases from 220 meV at 50k samples to 41 meV at 800k samples.
+Similarly, the force MAE decreases from 241 meV/Åat 5k samples to 54 meV/Åat 80k samples.
+Furthermore, deeper neural networks improve performance faster with the growth of the dataset size.
+This trend highlights the scalability of AlphaNet with increasing dataset size and model size.
+(a) Energy MAE (meV) vs Data Size
+(b) Force MAE (meV/Å) vs Data Size
+Figure 2: Mean absolute errors of energy and force predictions of models with $L$ number of layers as a function of training data size on the zeolite dataset.
+2.3 AlphaNet is efficient—achieves the best speed-accuracy trade-off
+The primary advantage of NNIP models over density functional theory (DFT) calculated energy and force lies in their significantly faster inference speed.
+Although increasing the number of parameters in an NNIP model generally enhances its performance, this often comes at the cost of slower inference speed.
+To evaluate the speed of different NNIPs, we evaluate the average forward pass time across 2,000 zeolite structures with a batch size of 10.
+In addition, we test systems of varying sizes by expanding the unit cells.
+Although we aim to maintain a consistent number of parameters across models, slight variations exist due to different components in each model (detailed in Section B.4).
+Figure 3: Comparisons of inference speed among various NNIP models.
 (a) Inference speed as a function of system size.
-The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches.
-(b) Inference speed and F1 performance of models accessible and compliant with the Matbench Discovery leaderboard.
-The speed is calculated as the number of energies predicted per second.
+The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches on the Zeolite dataset.
+(b) Inference speed, F1 performance and model size of models accessible and compliant with the Matbench Discovery leaderboard.
+The speed is calculated as the number of energy evaluation per second.
+As illustrated in Figure 3 (a), AlphaNet demonstrates a distinct speed advantage over other NNIP models, exhibiting a relatively slow increase in computational time as the size of the system increases.
+This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
+Furthermore, we evaluated the inference speed of the pre-trained models on the Matbench Discovery leaderboard (Riebesell et al., 2024).
+To ensure fairness, we select models that are compliant, open-source, and accessible without additional requests.
+We evaluate the average speed of energy prediction and present in Figure 3 (b).
+The models included AlphaNet, ORB-v2 (Neumann et al., 2024), SevenNet (Park et al., 2024), GRACE (Bochkarev et al., 2024), MACE (Batatia et al., 2022), and CHGNet (Deng et al., 2023).
+Among these models, GRACE, developed based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based.
+Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-l3i5).
+This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.
+Table 5: Efficiency evaluated on the OC20 dataset, using a single A100 40GB GPU. Conserv. denotes energy conservation. Boldface indicates the best performance.
+Models Parameters $\downarrow$ Training Speed (Sample/Sec) $\uparrow$ Training Memory (GB/Sample) $\downarrow$ Inference Speed (Sample/Sec) $\uparrow$ Inference Memory (GB/Sample) $\downarrow$ Conserv.
+SchNet 9.1M 39 2.5 112.1 1.7 DimeNet++ 10.1M 29 3.6 92 2.5 GemNet-OC 38M 16.22 1.63 39.2 1.61 $\times$ eSCN 51M 8.5 1.75 25.2 1.74 $\times$ EScAIP 83M 37.8 1.23 111.8 1.09 $\times$ EquiformerV2 31M 15.7 1.63 36.72 1.61 $\times$ AlphaNet 6.1M 41.2 2.2 120.1 1.5 $\checkmark$
+The last efficiency evaluation is on the OC20 dataset, as shown in Table 5.
+Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than ut.
+It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for PyTorch autograd, resulting in higher inference time and memory costs.
 3 Discussion
-AlphaNet presents a significant advancement as a scalable machine learning potential, offering both high precision and fast computational speed.
+AlphaNet presents a significant advancement as a scalable neural network interatomic potential, offering both high precision and fast computational speed.
 This success can be attributed to its local-frame-based architecture, which scalarizes geometric quantities over local frames and effectively aggregate local frames to construct local- and global-aware features.
@@
 One of the key advantages of AlphaNet lies in its scalability.
-Specifically, our model is capable of maintaining high performance as the size of the dataset increases, which sets it apart from other large-scale atomistic models.
+Specifically, our model is capable of maintaining high performance as both the size of the dataset and size of the atomic systems increase.
 Despite the growing complexity introduced by larger datasets, AlphaNet operates with fewer parameters, making it more efficient while still delivering robust results.
 This ability to handle large datasets without sacrificing accuracy positions AlphaNet as a promising solution for extensive simulations and real-world applications.
-In future work, we aim to extend the model’s capabilities to explore additional types of interactions, such as hydrogen bonding, which are critical in a wide range of chemical and biological systems.
-Another avenue for further exploration is the interpretability of the model.
+For future work, we aim to extend the model’s capabilities to explore additional types of interactions, such as hydrogen bonding, which are critical in a wide range of chemical and biological systems.
+Furthermore, the interpretability of the proposed model is worth exploring.
 While our current implementation achieves impressive results, gaining a deeper understanding of how the model interprets and processes interactions at the atomic level will enhance its applicability and trustworthiness in practical scenarios.
-Despite these advancements, one challenge remains: the efficient use of GPU memory during training, particularly when dealing with large systems consisting of over 1,000 atoms (though we can parallelize it across multiple GPUs).
-This is a common limitation shared by most message-passing neural networks (MPNNs), and addressing it will be crucial for enabling the model to scale further in both size and complexity.
+There is another long-term challenge: the efficient use of GPU memory during training, particularly when dealing with large systems consisting of over 1,000 atoms (though we can parallelize it across multiple GPUs).
+This is a common limitation shared by most message-passing neural networks, and addressing it will be crucial for enabling the model to scale further in both size and complexity.
 Future work will focus on optimizing memory usage to overcome this bottleneck and ensure the model’s applicability to even larger systems.
-Improving the smoothness of the output is also crucial, as we believe it would enhance the models’ generalization ability and yield a more physically meaningful potential energy surface.
 4 Methodology
-4.1 Problem Formulation
-Machine learning forcefields.
+Neural Network Interatomic Potentials.
 A single state of a molecular system can be described by a set of atom types $h\in\mathbb{R}^{n\times d}$ and atomic positions $x\in\mathbb{R}^{n\times 3}$.
 In molecular dynamics, we are interested in learning a function $f:\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}$ which predicts the energy of the current state of the molecular system readily to be used for simulation.
-The force can be further calculated by $-\nabla_{x}f(x)$, thus often referred as to machine learning force fields (MLFFs).
+The force can be further calculated by $-\nabla_{x}f(x)$, thus often referred as to neural network interatomic potentials (NNIPs).
 Message-passing neural networks.
@@
 In addition to permutation, we need to consider the special Euclidean group in 3D, SE(3), which includes rotation and translation operations.
-We do so by leveraging local frames [11].
+We do so by leveraging local frames (Du et al., 2022).
 Specifically, we build a set of equivariant and complete frames and scalarize the geometric quantity which does not break SE(3)-equivariance under nonlinear neural network encodings.
-To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following [11]
+To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following (Du et al., 2022):
 $\mathcal{F}_{ij}=\left(\hat{e}^{1}_{ij},\hat{e}^{2}_{ij},\hat{e}^{3}_{ij}\right)=\left(\frac{x_{i}-x_{j}}{\|x_{i}-x_{j}\|},\ \frac{(x_{i}-\overline{x})\times(x_{j}-\overline{x})}{\|(x_{i}-\overline{x})\times(x_{j}-\overline{x})\|},\ \frac{(x_{i}-x_{j})\times\left((x_{i}-\overline{x})\times(x_{j}-\overline{x})\right)}{\|(x_{i}-x_{j})\times\left((x_{i}-\overline{x})\times(x_{j}-\overline{x})\right)\|}\right)$
-To scalarize geometric quantity such as a vector $v$, we take the inner product between it and the frames which results in a set of invariant scalars and can be transformed back through a tensorization operation
+To scalarize geometric quantity such as a vector $v$, we take the inner product between it and the frames which results in a set of invariant scalars and can be inverted via a tensorization operation
 $\text{Scalarize}(v,\mathcal{F}_{ij})=(s^{1}_{ij},s^{2}_{ij},s^{3}_{ij})=(e^{1}_{ij}\cdot v,e^{2}_{ij}\cdot v,e^{3}_{ij}\cdot v)$
@@
 Efficient and expressive equivariant message-passing neural networks.
-Following [12], we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames.
+Following (Du et al., 2024), we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames.
 We consider two modules to improve the expressiveness while maintaining the efficiency of the network.
@@
 Rotary positional embedding as invariant frame transition.
-Positional embedding are known to improve and stabilize the training of transformers [34].
-We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture [37].
+Positional embedding are known to improve and stabilize the training of transformers (Shaw et al., 2018).
+We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture (Su et al., 2024).
 In fact the frame transition can be considered as a generalized rotary positional embedding for 3D equivariant frames.
 Note that the spirit of rotary position embedding is to design an embedding scheme for relative position in sequential data.
-To build the rotary positional embedding,  [37] utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
+To build the rotary positional embedding,  (Su et al., 2024) utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
 $\langle f_{q}(x_{i},i),f_{k}(x_{j},j)\rangle=g(x_{i},x_{j},i-j)$
 where $f_{q}$ and $f_{k}$ denote the neural network embeddings of query and key inside an attention layer.
-Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices [37]:
+Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices (Su et al., 2024):
 $\begin{pmatrix}\cos\theta_{n},&-\sin\theta_{n}\\ \sin\theta_{n},&\cos\theta_{n}\end{pmatrix}$
@@
 Temporal connection.
-In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network [7].
+In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network (Chen et al., 2018).
 As message-passing neural networks are local, we believe it is beneficial to integrate multi-scale information.
@@
 […]
+Supplementary Information
+AlphaNet: Scaling Up Local Frame-based NNIP
+Appendix A Local Structures, Completeness, and Expressive Power of GNNs
+Graph neural networks (GNNs) derive their locality from aggregating information within $k$-hop neighborhoods, typically limited to 1-hop interactions.
+While this design enables efficient learning, it inherently restricts their focus to local substructures such as trees, triangles, or cycles.
+These substructures define distinct local isomorphism types: (1) Tree isomorphism, where subgraphs share identical hierarchical parent-child relationships; (2) Triangular isomorphism, requiring matching cyclic triplets (e.g., molecular bond angles); and (3) Subgraph isomorphism, preserving node features and connectivity patterns.
+The expressive power of permutation-invariant GNNs is classically evaluated via the 1-Weisfeiler-Lehman (1-WL) test  (Xu et al., 2019).
+As established by  (Wijesinghe and Wang, 2022), the 1-WL test is equivalent to a GNN’s capacity to discriminate local subtree isomorphism, measuring its ability to distinguish hierarchical neighborhood structures.
+However, this framework operates solely on 2D graph topologies, ignoring 3D geometric relationships (e.g., bond angles, torsional dihedrals), thereby failing to address molecular or material systems where spatial configurations are decisive.
+Completeness in GNNs requires a transformation function $f$ to satisfy $f(X)=f(Y)\iff X\text{ and }Y\text{ are isomorphic}$.
+We can extends completeness to local geometric isomorphism, where isomorphism is defined under 3D geometric constraints (distances, angles, chiralities).
+Concurrently, methods encoding 3D geometric invariants  (Wijesinghe and Wang, 2022) have demonstrated enhanced expressivity, enabling differentiation of graphs with distinct local atomic environments beyond WL limitations.
+Similar to LeftNet (Du et al., 2023), AlphaNet is able to differentiate local structures beyond tree isomorphism by defining the 3D structure weights.
+Appendix B Experimental Details
+B.1 Datasets
+The Defected Bilayer Graphene Dataset (Ying et al., 2024) consists of reference structures specifically designed to train and validate machine learning potentials (MLPs).
+It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers).
+The dataset comprises single-point DFT (PBE+MBD) energies and atomic forces calculated for various configurations, including different interlayer distances, stacking modes, and manually deformed structures.
+Additionally, snapshot configurations from MD simulations at different temperatures were included.
+The data were split into training, validation, and test sets, containing 3988, 4467, and 200 structures, respectively.
+Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.
+The Formate Decomposition on Cu Dataset (Batzner et al., 2022) consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond.
+The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase.
+Nudged elastic band (NEB) method (Henkelman et al., 2000) was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations.
+These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory.
+AlphaNet was trained on 2500 uniformly sampled structures from the full dataset, with a validation set of 250 structures and the mean absolute error evaluated on the remaining structures.
+The Zeolite Dataset comprises essential reference structures designed for various applications such as catalysis, adsorption, and separation.
+Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability.
+The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) (Car and Parrinello, 1985) simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) (Kresse and Furthmüller, 1996; Kresse and Hafner, 1993).
+For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration.
+The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
+The OC2M Dataset (Chanussot et al., 2021) is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials.
+This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both).
+Each of them contains around 1 million structures.
+It encompasses configurations involving 56 different chemical elements, ensuring a broad and diverse representation of atomic environments.
+These configurations were generated using density functional theory (DFT) calculations, with a focus on a wide range of materials science applications, including bulk materials, surfaces, and defect structures.
+Originating from the comprehensive OC20 Dataset, the OC2M subset provides accurate total energies and atomic forces for each configuration.
+This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
+The MPTraj Dataset and Matbench Discovery Test (Riebesell et al., 2024).
+MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials (Deng et al., 2023).
+Matbench Discovery is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set (Wang et al., 2021).
+In this case we train our model on the MPTraj dataset and test it on the WBM test set.
+B.2 Hyperparamters
+Table 6: Hyperparameter configurations for five different datasets.
+Gra.
+denotes the defected graphene dataset, For.
+denotes the formate decomposition dataset, Zeo.
+denotes the zeolite dataset, OC2M denotes the OC2M dataset, MPTraj denotes the MPTraj dataset.
+Parameter Gra. For. Zeo. OC2M MPTraj
+Learning rate (lr) 5e-4 5e-4 5e-4 1e-4 1e-4
+Scheduler Cos. Cos. Cos. Cos. Cos.
+Epochs 150 150 100 130 200
+Loss func. MAE MAE MAE MAE MAE
+Batch size (per device) 24 12 96 4 12
+Num layers 3 3 3 4 6
+Heads 16 16 16 16 24
+Num basis 32 32 32 96 256
+Hidden channels 128 128 128 256 256
+Cutoff (Å) 5 5 5 5 6
+B.3 Training cost on the OC2M dataset and MPTraj
+Table 7: Training cost for the OC2M and MPTraj datasets. The training memory here is determined by dividing the fixed memory capacity of a GPU by the batch size.
+Dataset Training Time (GPU Days) Number of Parameters Training Memory (GB per sample)
+OC2M 361 6.1M 5.1
+MPTraj 243 16.2M 10.1
+B.4 Baseline Models
+For computational modeling configurations, NequIP achieves optimal performance with tensor rank 2, feature size 32, and 5 Åcutoff in the formate decomposition dataset, while using tensor rank 1 with 7 Åcutoff for defective graphene studies.
+The zeolite dataset employs a fine-tuned DPA-2 potential via Deep Pot implementation ((Zhang et al., 2024b)).
+Models trained on OC20 leverage pre-trained SchNet/DimeNet++ architectures from https://fair-chem.github.io/core/model_checkpoints.html, with other results cross-validated against (Qu and Krishnapriyan, 2024) using publicly accessible validation sets.
+MPtrj benchmark data integrates results from https://matbench-discovery.materialsproject.org/ and comparative analysis in (Qu and Krishnapriyan, 2024).
+Speed testing on zeolite structures maintains uniform 2-layer architectures with 5 Åcutoff, where PaiNN/SchNet employ 120-atom basis and 16 RBF basis, MACE uses tensor order 0-1 (feature size 16), NequIP extends to tensor order 0-2 (feature size 16), AlphaNet configures 32 hidden channels with 12 attention heads, and EquiformerV2 implements sphere/edge channels 16 with 16-dimensional attention hidden states and 2-head mechanisms.
+Appendix C Ablation Study
+Table 8: Ablation study on the impact of RoPE, temporal connections (Temp Connect), and frame transition encoding (FTE).
+Configuration Energy (meV/atom) $\downarrow$ Force MAE (meV/Å) $\downarrow$
+Full Model 1.7 32.0
+w/o RoPE 25.9 113.6
+w/o Temp Connect 2.9 38.2
+w/o FTE 2.8 45.3
+We do ablation study on the Defected Graphene dataset,  Table 8 shows that all the component matters, but it is worth notice that our model can not get down the loss and performs bad with out RoPE, though this did not happen in our other tests.
+While shows that RoPE is important, some other aspects may need to be considered such as more efficient way of optimizing tensor network.
+Appendix D Data Scaling and Dynamic loss weights
+The experimental datasets were generated using VASP (Kresse and Furthmüller, 1996; Kresse and Hafner, 1993), where the energy per atom typically converges around $5\,\mathrm{eV}$ under our default training settings.
+These configurations demonstrated robust performance within this energy regime.
+However, when applied to systems with significantly higher energy magnitudes per atom and greater energy variance—such as those produced by CP2K (Kühne et al., 2020) simulations—the original settings failed to achieve stable convergence.
+To address this challenge, we developed a dynamic optimization strategy: initially adjusting the energy-to-force loss weight ratio from $\lambda_{E}:\lambda_{F}=4:100$ to $0.05:100$, then progressively increasing $\lambda_{E}$ while simultaneously decaying the learning rate.
+This coordinated adaptation proved critical for successful training in high-variance scenarios.

### v4

@@
-AlphaNet: Scaling Up Local Frame-based Neural Network Interatomic Potentials
+AlphaNet: Scaling Up Local-frame-based Neural Network Interatomic Potentials
 Abstract
-Molecular dynamics simulations demand unprecedented accuracy and scalability to tackle grand challenges in energy materials, catalytic processes, and biomolecular design.
-To bridge this gap, we present AlphaNet, a local frame-based equivariant model that simultaneously advances computational efficiency and predictive precision for atomistic systems.
+Molecular dynamics simulations demand an unprecedented combination of accuracy and scalability to tackle grand challenges in catalysis and materials design.
+To bridge this gap, we present AlphaNet, a local-frame-based equivariant model that simultaneously improves computational efficiency and predictive precision for interatomic interactions.
 By constructing equivariant local frames with learnable geometric transitions, AlphaNet encodes atomic environments with enhanced representational capacity, achieving state-of-the-art accuracy in energy and force predictions.
-Extensive benchmarks—spanning defected graphene, formate decomposition, inorganic bulks, and large-scale datasets (OC2M and Matbench Discovery) demonstrate its superior performance over existing neural network interatomic potentials while ensuring scalability across diverse system sizes.
-The synergy of accuracy, efficiency, and transferability positions AlphaNet as a transformative tool for simulating multiscale phenomena, from catalyst dynamics to energy storage interfaces, with direct implications for accelerating the discovery of functional materials and complex molecular systems.
-Our code and data is available at https://github.com/zmyybc/AlphaNet.
-00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors.
-Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@gmail.com and haixiao@tsinghua.edu.cn.
+Extensive benchmarks on large-scale datasets spanning molecular reactions, crystal stability, and surface catalysis (Matbench Discovery and OC2M) demonstrate its superior performance over existing neural network interatomic potentials while ensuring scalability across diverse system sizes with varying types of interatomic interactions.
+The synergy of accuracy, efficiency, and transferability positions AlphaNet as a transformative tool for modeling multiscale phenomena, decoding dynamics in catalysis and functional interfaces, with direct implications for accelerating the discovery of complex molecular systems and functional materials.
+Our code and data are available at https://github.com/zmyybc/AlphaNet.
 1 Introduction
-Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems (Richards et al., 2016; Boero et al., 1998; Lindorff-Larsen et al., 2011).
-While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps).
-As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
-Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories (Jones and Chapman, 1924; Cornell et al., 1995; MacKerell et al., 1998).
-However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics (Iftimie et al., 2005).
+Molecular dynamics (MD) simulations have become essential for investigating and elucidating complex phenomena across diverse fields, including biology, catalysis, and energy engineering [21, 20, 1, 2, 27, 43, 9, 37].
+While atomic forces necessary for MD simulations can be derived from quantum mechanical approaches such as density functional theory (DFT) [10], the computational cost associated with such first-principles methods severely limits their applications to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps).
+Consequently, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
+Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories [29, 14, 38, 47, 55, 31].
+However, the simplicity of these models often compromises their accuracy, necessitating a delicate trade-off between computational efficiency and the fidelity of the simulated dynamics, especially for reactive chemistry [55, 53, 28, 64].
 The advent of machine learning techniques has introduced a promising solution to this dilemma.
-By training models on data derived from first-principles calculations, neural network interatomic potentials (NNIPs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields.
-Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, NNIPs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
+By training models on data derived from first-principles calculations, neural network interatomic potentials (NNIPs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields [4, 5, 63, 6, 3, 51, 32].
+Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, NNIPs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities [7].
 As interest in applying NNIPs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models.
-One example among the others, DPA-1 (Zhang et al., 2024a), is a large-scale pre-trained model with improved attention architecture over DeepPot (Wang et al., 2018).
-Similarly, JMP (Shoghi et al., 2024) leverages diverse molecular systems of different types as a joint pre-training model.
-MACE (Batatia et al., 2022), on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions.
-Collectively, these models represent significant strides in the evolution of NNIP-driven atomistic simulations (Qu and Krishnapriyan, 2024; Liao et al., 2024; Gasteiger et al., 2021; Batzner et al., 2022; Schütt et al., 2018; Smith et al., 2017; Deng et al., 2023).
-Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models.
+One example among the others, DPA-1 [62], is a large-scale pre-trained model with improved attention architecture over DeepPot [56].
+Similarly, JMP [49] leverages diverse molecular systems of different types as a joint pre-training model.
+MACE [4], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions.
+Collectively, these models represent significant strides in the evolution of NNIP-driven atomistic simulations [42, 36, 23, 5, 46, 50, 15, 13, 60].
+Nevertheless, the balance between computational efficiency and accuracy is key to practical applications of these models.
 On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy.
@@
 Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection.
-Thus, equivariant models are developed to respect these symmetries  (Satorras et al., 2021; Thomas et al., 2018), and most of these expressive models by date (Batzner et al., 2022; Batatia et al., 2022; Liao et al., 2024) are based on the spherical harmonics (Geiger and Smidt, 2022), but calculating the tensor product of irreducible representations imposes expensive computational overheads.
-While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global (Jumper et al., 2021; Du et al., 2022, 2023; Duval et al., 2023), and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
-In this paper, we propose a local-frame-based equivariant atomistic model, named AlphaNet, for accurate yet highly efficient atomistic simulations.
-Building on the success of frame-based NNIPs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling.
-Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art (SOTA) NNIP models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, surface reactions and inorganic bulks.
+Thus, equivariant models are developed to respect these symmetries  [45, 54], and most of these expressive models by date [5, 4, 36] are based on the spherical harmonics [25], but calculating the tensor product of irreducible representations imposes expensive computational overheads.
+While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global [30, 16, 17, 19], and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
+Herein, we propose a local-frame-based equivariant atomistic model, named AlphaNet, for accurate yet highly efficient atomistic simulations.
+Building on the success of frame-based NNIPs in small atomistic systems, we introduce an additional rotary position embedding to enhance the frame transition and temporal connection for multi-scale modeling.
+Extensive quantitative benchmarks demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art (SOTA) NNIP models on a variety of datasets, including formate decomposition, defected graphenes, zeolites, Open Catalyst 2020 (OC20), and Matbench discovery WBM test set, which cover various types of interatomic interactions across a rich set of systems and phenomena, ranging from molecules and bulks to surfaces, from gas-phase reactions and crystal stability to surface catalysis.
 2 Results
-We systematically assess the capabilities of AlphaNet by examining its prediction accuracy, scalability with respect to model and dataset size, and computational efficiency in terms of inference speed and memory usage.
+We systematically assess the capabilities of AlphaNet by examining its prediction accuracy, scalability with respect to model and dataset sizes, as well as computational efficiency in terms of inference speed and memory usage.
 Initially, we validate atomic-level prediction accuracy across five diverse chemical systems, simultaneously demonstrating AlphaNet’s scalability concerning dataset and system sizes.
@@
 Comprehensive details regarding datasets, comparative methodologies, and hyperparameter configurations are provided in Appendix B.
-2.1 AlphaNet achieving high accuracy across all datasets.
+2.1 AlphaNet achieving high accuracy across all datasets
 We first consider the formate decomposition dataset, which serves as a representative example of catalytic surface reactions, specifically focuses on the dehydrogenation of formate (HCOO* → H* + CO2) on a Cu $\langle 110\rangle$ surface.
-For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Åfor force and 0.23 meV/atom for energy, compared to NequIP (Batzner et al., 2022) 47.3 meV/Åand 0.50 meV/atom, respectively, as shown in Table 1.
+For this dataset, AlphaNet achieves a mean absolute error (MAE) of 45.5 meV/Åfor force and 0.23 meV/atom for energy, compared to NequIP [5] 47.3 meV/Åand 0.50 meV/atom, respectively, as shown in Table 1.
 These results underscore AlphaNet’s ability to effectively capture the intricate nature of heterogeneous systems, which often involve both metallic and covalent bonding, along with complex charge transfer processes between the metal and adsorbed molecules.
 This high degree of accuracy highlights the model’s suitability for simulating catalytic reactions with multiple interaction types.
-Another challenge for NNIP models is accurately modeling the sliding layer effect in materials like defected graphene, where inter-layer interactions are critical.
-On the defected graphene dataset, AlphaNet achieves a force MAE of 32.0 meV/Åand an energy MAE of 1.7 meV/atom, substantially outperforming NequIP, which records 60.2 meV/Åfor force MAE and 1.9 meV/atom for energy MAE.
-Furthermore, AlphaNet successfully reproduces the binding energy curve for AB-stacked bilayer graphene as calculated by PBE+MBD, without requiring explicit long-range dispersion interaction corrections.
-The deviation in the shallow sliding potential energy landscape is less than 0.4 meV/atom, illustrating AlphaNet’s capacity to handle systems with subtle interlayer forces and complex structural dynamics (as shown in Figure 1).
 Table 1: Results on the Formate Decomposition and Defected Graphene datasets. Boldface indicates the best performance.
@@
 Energy MAE (meV/atom) $\downarrow$ 0.50 0.23
-Defected Graphene Force MAE (meV/Å) $\downarrow$ 60.2 32.0
-Energy MAE (meV/atom) $\downarrow$ 1.9 1.7
+Defected Graphene Force MAE (meV/Å) $\downarrow$ 60.2 19.4
+Energy MAE (meV/atom) $\downarrow$ 1.9 1.2
+A further challenge for neural network interatomic potentials (NNIPs) is the accurate modeling of inter-layer sliding effects in layered materials such as defected graphene, where interlayer interactions are essential.
+On the defected graphene dataset, AlphaNet attains a force MAE of 32.0 meV/Å and an energy MAE of 1.7 meV/atom, significantly outperforming NequIP’s 60.2 meV/Å and 1.9 meV/atom, respectively.
+Additionally, AlphaNet successfully reproduces the binding energy profile for AB-stacked bilayer graphene, benchmarked against PBE+MBD calculations, without explicitly incorporating long-range dispersion corrections.
+Deviations within the shallow sliding potential energy landscape are less than 0.4 meV/atom, highlighting AlphaNet’s exceptional performance in modeling subtle interlayer forces and complex structural dynamics (see Figure 1).
 (a) Parity plot for total energies obtained for the training (blue), validation (orange), and test (green) data sets.
@@
 Figure 1: Comparisons between AlphaNet predictions and DFT calculations on the Zeolite dataset.
-The third dataset we consider is the zeolite dataset, comprising 16 types of zeolites and a total of 800,000 configurations.
-Table 2 shows the results of Deep Pot (Wang et al., 2018) and AlphaNet across different zeolite configurations.
-Across nearly all tested configurations, AlphaNet significantly outperforms Deep Pot in both energy and force predictions.
-For instance, in the ABWopt configuration, AlphaNet achieves an Energy MAE of 41 meV, substantially lower than Deep Pot’s 90 meV.
-Similarly, AlphaNet records a Force MAE of 54 meV/Å, outperforming Deep Pot’s 90 meV/Å.
+The third dataset evaluated is the zeolite dataset, consisting of 16 distinct zeolite types and comprising a total of 800,000 configurations.
+As summarized in Table 2, AlphaNet demonstrates significantly improved performance compared to Deep Pot [56] across nearly all zeolite configurations evaluated, exhibiting superior accuracy in both energy and force predictions.
+For example, in the ABWopt configuration, AlphaNet achieves an energy mean absolute error (MAE) of 12 meV, markedly lower than Deep Pot’s 90 meV.
+Similarly, AlphaNet attains a force MAE of 19 meV/Å, clearly outperforming the 90 meV/Å obtained by Deep Pot.
 Table 2: Results on the Zeolite dataset. Boldface indicates the best performance. (Unit for energy is meV/atom and unit for force is meV/Å).
@@
 Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
-Energy MAE $\downarrow$ 90 71 110 73 210 41 150 62
-Force MAE $\downarrow$ 90 54 50 55 60 90 90 10
+Energy MAE $\downarrow$ 90 12 110 6.8 210 29 150 18
+Force MAE $\downarrow$ 90 19 50 12 60 16 90 15
 EDIopt FERopt GISopt JBWopt
 Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
-Energy MAE $\downarrow$ 40 5.4 290 100 60 56 150 18
-Force MAE $\downarrow$ 50 59 130 18 50 3.1 70 50
+Energy MAE $\downarrow$ 40 8.7 290 43 60 11 150 10
+Force MAE $\downarrow$ 50 13.5 130 28 50 11 70 13
 LOSopt LTAopt LTJopt NATopt
 Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
-Energy MAE $\downarrow$ 110 130 150 90 70 1.2 210 47
-Force MAE $\downarrow$ 70 46 64 60 110 60 110 9.9
+Energy MAE $\downarrow$ 110 21 150 19 70 15 210 16
+Force MAE $\downarrow$ 70 12 64 12 110 11 110 15
 PARopt PHIopt SODopt THOopt
 Metric Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet Deep Pot AlphaNet
-Energy MAE $\downarrow$ 90 6.5 60 16 200 9.5 160 70
-Force MAE $\downarrow$ 70 3.5 120 50 110 21 60 47
+Energy MAE $\downarrow$ 90 18 60 20 200 9.5 160 17
+Force MAE $\downarrow$ 70 24 120 14 110 12 60 15
 Table 3: Results on the OC20 validation set.
@@
 AlphaNet ($\lambda_{E}=20$) 6.1M 0.24 0.062
-The next challenge, we consider the Open Catalyst Project OC20 dataset (Chanussot et al., 2021), a comprehensive collection of crystal structures for training NNIP models in catalysis.
-We focus on the subset OC2M and the Structure-to-energy-and-force (S2EF) task.
-We compare the performance of our model against established benchmarks, including EquiformerV2 (Liao et al., 2024), EScAIP (Qu and Krishnapriyan, 2024), eSCN (Passaro and Zitnick, 2023), GemNet-OC (Gasteiger et al., 2022), SchNet (Schütt et al., 2018), and DimeNet++ (Gasteiger et al., 2020).
-AlphaNet, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV.
-This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit worse MAE values (>0.35 eV).
-However, AlphaNet trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP in the force prediction task.
-This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$.
+In addressing the next challenge, we consider the Open Catalyst Project OC20 dataset [11], a comprehensive collection of crystal structures specifically curated for training neural network interatomic potential (NNIP) models in catalysis applications.
+Our study focuses on the OC2M subset and evaluates performance on the Structure-to-Energy-and-Force (S2EF) task.
+We benchmark our proposed model, AlphaNet, against established state-of-the-art methods, including EquiformerV2 [36], EScAIP [42], eSCN [41], GemNet-OC [24], SchNet [46], and DimeNet++ [22].
+AlphaNet, trained on the 2M subset for 2,000,000 steps, achieves remarkable accuracy in energy prediction, yielding a mean absolute error (MAE) of 0.24 eV.
+This accuracy is on par with larger-scale models such as EquiformerV2 and EScAIP.
+In contrast, models like SchNet and DimeNet++, even when trained on the full dataset, demonstrate inferior performance, with MAE values exceeding 0.35 eV.
+However, AlphaNet, trained exclusively on the smaller 2M subset, does not surpass the larger-scale models (EquiformerV2 and EScAIP) in force prediction tasks.
+This discrepancy likely arises from the inherent conservativeness of the model architectures, which enforce consistency by computing forces strictly as negative gradients of the predicted energies ($-\nabla_{x}E(x)$).
 Table 4: Results on the Matbench Discovery WBM unique prototype dataset. Boldface indicates the best performance.
@@
 eqV2 S DeNS 0.815 5.042 0.771 0.941 0.036 0.788 31.2M EFSD
-EScAIP-Small 0.782 5.634 0.712 0.939 0.038 0.783 45M EFSD
-AlphaNet 0.799 4.863 0.743 0.933 0.041 0.745 16.2M EFSG
-For the last challenge, we validate the effectiveness of AlphaNet on the well-established Matbench discovery WBM test set (Riebesell et al., 2024), AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4.
-With a 0.799 F1 score and 4.863 DAF score, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints.
-Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$).
-This combination of efficiency and physics-aware design positions AlphaNet as a robust NNIP model for scalable materials modeling.
-2.2 AlphaNet is scalable—effective scaling with both model, data and system sizes
-In addition to maintaining high performance with large datasets and large systems, we validate the scaling of model size and dataset size on the zeolite dataset.
-We can observe the performance improves as the training data size increases from 50k to 800k samples, as shown in Figure 2.
-Specifically, the energy MAE decreases from 220 meV at 50k samples to 41 meV at 800k samples.
-Similarly, the force MAE decreases from 241 meV/Åat 5k samples to 54 meV/Åat 80k samples.
-Furthermore, deeper neural networks improve performance faster with the growth of the dataset size.
-This trend highlights the scalability of AlphaNet with increasing dataset size and model size.
+EScAIP 0.782 5.634 0.712 0.939 0.038 0.783 45M EFSD
+AlphaNet-L 0.799 4.863 0.743 0.933 0.041 0.745 16.2M EFSG
+AlphaNet-S*111* means not submitted yet 0.808 4.915 0.751 0.935 0.037 0.796 4.5M EFSG
+Lastly, we validate the effectiveness of AlphaNet using the established Matbench discovery WBM test set [44].
+AlphaNet demonstrates an excellent balance between predictive accuracy and physical consistency on the WBM benchmark (Table 4).
+Achieving an F1 score of 0.799 and a DAF score of 4.863, AlphaNet attains near state-of-the-art classification performance while utilizing an EFSG-type conservative force field to rigorously enforce physical constraints.
+Notably, despite its compact architecture (16.2M parameters), AlphaNet-L surpasses the larger eqV2 S DeNS model (31.2M parameters) in regression tasks, yielding an improved coefficient of determination ($R^{2}=0.742$) and lower prediction error (mean absolute error, $\text{MAE}=0.039$).
+There is also a new small size model(AlphaNet-S) with a different RBF function which performs even better.
+This exceptional combination of computational efficiency and physics-informed design positions AlphaNet as a powerful neural network interatomic potential (NNIP) model, ideally suited for scalable materials modeling.
+2.2 AlphaNet scaling effectively with model, data and system sizes
+In addition to demonstrating robust performance on large datasets and extensive systems, we systematically evaluate how model size and dataset scale influence performance using the zeolite dataset.
+As depicted in Figure 2, increasing the training dataset size from 50k to 800k samples notably improves prediction accuracy.
+Specifically, the mean absolute error (MAE) for energy predictions decreases significantly from 220 meV at 50k samples to 41 meV at 800k samples.
+Similarly, the force prediction MAE reduces markedly from 241 meV/Å at 5k samples to 54 meV/Å at 80k samples.
+Furthermore, deeper neural networks exhibit accelerated performance improvements as dataset size increases, underscoring AlphaNet’s scalability with respect to both dataset and model size.
 (a) Energy MAE (meV) vs Data Size
@@
 Figure 2: Mean absolute errors of energy and force predictions of models with $L$ number of layers as a function of training data size on the zeolite dataset.
-2.3 AlphaNet is efficient—achieves the best speed-accuracy trade-off
-The primary advantage of NNIP models over density functional theory (DFT) calculated energy and force lies in their significantly faster inference speed.
-Although increasing the number of parameters in an NNIP model generally enhances its performance, this often comes at the cost of slower inference speed.
-To evaluate the speed of different NNIPs, we evaluate the average forward pass time across 2,000 zeolite structures with a batch size of 10.
-In addition, we test systems of varying sizes by expanding the unit cells.
-Although we aim to maintain a consistent number of parameters across models, slight variations exist due to different components in each model (detailed in Section B.4).
+2.3 AlphaNet demonstrating enhanced speed-accuracy trade-off
+The primary advantage of NNIP models over density functional theory (DFT)-calculated energies and forces is their substantially faster inference speed.
+However, increasing the number of parameters in an NNIP model to enhance its predictive accuracy typically results in slower inference speeds.
+To systematically assess inference speed across various NNIPs, we measure the average forward-pass computation time using 2,000 zeolite structures, processed with a batch size of 10.
+Additionally, we evaluate inference performance across systems of varying sizes by systematically expanding their unit cells.
+While we strive to maintain a consistent parameter count across models, minor variations persist due to differing internal architectures and components, as detailed in Section B.4.
+(a) Inference speed as a function of system size on the zeolite dataset.
+(b) Inference speed vs performance on Matbench Discovery.
 Figure 3: Comparisons of inference speed among various NNIP models.
-(a) Inference speed as a function of system size.
+(a).
 The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches on the Zeolite dataset.
-(b) Inference speed, F1 performance and model size of models accessible and compliant with the Matbench Discovery leaderboard.
+(b).
+All models are accessible and compliant with the Matbench Discovery leaderboard.
 The speed is calculated as the number of energy evaluation per second.
-As illustrated in Figure 3 (a), AlphaNet demonstrates a distinct speed advantage over other NNIP models, exhibiting a relatively slow increase in computational time as the size of the system increases.
-This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
-Furthermore, we evaluated the inference speed of the pre-trained models on the Matbench Discovery leaderboard (Riebesell et al., 2024).
-To ensure fairness, we select models that are compliant, open-source, and accessible without additional requests.
-We evaluate the average speed of energy prediction and present in Figure 3 (b).
-The models included AlphaNet, ORB-v2 (Neumann et al., 2024), SevenNet (Park et al., 2024), GRACE (Bochkarev et al., 2024), MACE (Batatia et al., 2022), and CHGNet (Deng et al., 2023).
-Among these models, GRACE, developed based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based.
-Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-l3i5).
-This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.
+As shown in Figure 3(a), AlphaNet exhibits a clear computational advantage over other NNIP models, characterized by a notably slower increase in computational time with growing system size.
+This efficient scaling behavior makes AlphaNet particularly well-suited for simulations involving larger systems.
+Furthermore, we assess the inference speed of pre-trained models featured on the Matbench Discovery leaderboard [44].
+To ensure a fair comparison, we select compliant, open-source models that are freely accessible without additional authorization.
+The average energy prediction speeds of these models are presented in Figure 3(b).
+The evaluated models include AlphaNet, ORB-v2 [39], SevenNet [40], GRACE [8], MACE [4], and CHGNet [15].
+Among these, GRACE implemented using TensorFlow demonstrates significantly faster inference compared to the other models, which are built on PyTorch, and our model, AlphaNet-S, comes in the second place.
+In particular, AlphaNet-L exhibits an inference speed comparable to models possessing approximately ten times fewer parameters, such as SevenNet-l3i5.
+This finding indicates that AlphaNet maintains competitive computational efficiency despite its relatively larger size, raising an interesting prospect for investigating its potential performance with reduced parameters in future studies.
 Table 5: Efficiency evaluated on the OC20 dataset, using a single A100 40GB GPU. Conserv. denotes energy conservation. Boldface indicates the best performance.
@@
 SchNet 9.1M 39 2.5 112.1 1.7 DimeNet++ 10.1M 29 3.6 92 2.5 GemNet-OC 38M 16.22 1.63 39.2 1.61 $\times$ eSCN 51M 8.5 1.75 25.2 1.74 $\times$ EScAIP 83M 37.8 1.23 111.8 1.09 $\times$ EquiformerV2 31M 15.7 1.63 36.72 1.61 $\times$ AlphaNet 6.1M 41.2 2.2 120.1 1.5 $\checkmark$
-The last efficiency evaluation is on the OC20 dataset, as shown in Table 5.
-Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than ut.
-It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for PyTorch autograd, resulting in higher inference time and memory costs.
+The final efficiency evaluation is conducted using the OC20 dataset, with results summarized in Table 5.
+Notably, EScAIP demonstrates outstanding efficiency in both inference speed and memory usage, even given its substantial model size.
+AlphaNet exhibits a marginally faster inference speed; however, it incurs higher memory consumption compared to EScAIP.
+It is important to emphasize that models such as AlphaNet, SchNet, and DimeNet++ are conservative in their force calculations, inherently relying on PyTorch’s autograd functionality, thus necessitating additional computational overhead and consequently increasing inference time and memory usage.
 3 Discussion
-AlphaNet presents a significant advancement as a scalable neural network interatomic potential, offering both high precision and fast computational speed.
-This success can be attributed to its local-frame-based architecture, which scalarizes geometric quantities over local frames and effectively aggregate local frames to construct local- and global-aware features.
-By utilizing this architecture, our model excels at capturing complex interactions across a broad spectrum of systems, including metallic, covalent, and ionic bonding, as well as long-range interactions.
-One of the key advantages of AlphaNet lies in its scalability.
-Specifically, our model is capable of maintaining high performance as both the size of the dataset and size of the atomic systems increase.
-Despite the growing complexity introduced by larger datasets, AlphaNet operates with fewer parameters, making it more efficient while still delivering robust results.
-This ability to handle large datasets without sacrificing accuracy positions AlphaNet as a promising solution for extensive simulations and real-world applications.
-For future work, we aim to extend the model’s capabilities to explore additional types of interactions, such as hydrogen bonding, which are critical in a wide range of chemical and biological systems.
-Furthermore, the interpretability of the proposed model is worth exploring.
-While our current implementation achieves impressive results, gaining a deeper understanding of how the model interprets and processes interactions at the atomic level will enhance its applicability and trustworthiness in practical scenarios.
-There is another long-term challenge: the efficient use of GPU memory during training, particularly when dealing with large systems consisting of over 1,000 atoms (though we can parallelize it across multiple GPUs).
-This is a common limitation shared by most message-passing neural networks, and addressing it will be crucial for enabling the model to scale further in both size and complexity.
-Future work will focus on optimizing memory usage to overcome this bottleneck and ensure the model’s applicability to even larger systems.
+AlphaNet represents a significant advance in scalable neural network interatomic potentials, achieving exceptional accuracy while maintaining computational efficiency.
+Its success stems from a novel local-frame-based architecture that scalarizes geometric information within local frames and effectively aggregates these frames to produce features sensitive to both local and global chemical environments.
+Leveraging this architecture, AlphaNet adeptly captures complex atomic interactions across diverse chemical bonding types, including metallic, covalent, ionic, and long-range interactions, demonstrating versatility and broad applicability.
+One of AlphaNet’s key strengths is its exceptional scalability.
+The model consistently maintains high performance across varying dataset size and complexity of atomic systems.
+Remarkably, AlphaNet achieves robust accuracy while utilizing fewer parameters compared to conventional approaches, enhancing its computational efficiency.
+This capability to effectively manage large-scale datasets without compromising performance positions AlphaNet as a highly promising framework for extensive simulations and practical real-world applications.
+In future work, we plan to extend the model’s capabilities to account for additional interaction types, notably hydrogen bonding, which plays a critical role across diverse chemical and biological systems.
+Another important direction is enhancing the interpretability of the proposed model.
+Although our current implementation demonstrates strong predictive performance, its applicability and reliability in practical scenarios can be significantly improved by achieving deeper insight into how the model identifies, processes, and leverages atomic-level interactions.
+Additionally, addressing the efficient use of GPU memory during training presents a long-term challenge, particularly for large systems exceeding 1,000 atoms—even when leveraging parallelization across multiple GPUs.
+This limitation, common among most message-passing neural networks, must be overcome to facilitate scalability toward larger and more complex systems.
+Future research will prioritize memory optimization strategies to surmount this bottleneck, ensuring broader applicability and enabling analysis of increasingly intricate molecular systems.
 4 Methodology
@@
 In addition to permutation, we need to consider the special Euclidean group in 3D, SE(3), which includes rotation and translation operations.
-We do so by leveraging local frames (Du et al., 2022).
+We do so by leveraging local frames [16].
 Specifically, we build a set of equivariant and complete frames and scalarize the geometric quantity which does not break SE(3)-equivariance under nonlinear neural network encodings.
-To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following (Du et al., 2022):
+To build equivariant and complete frames from atomic positions, we consider the following edge-based frames following [16]:
 $\mathcal{F}_{ij}=\left(\hat{e}^{1}_{ij},\hat{e}^{2}_{ij},\hat{e}^{3}_{ij}\right)=\left(\frac{x_{i}-x_{j}}{\|x_{i}-x_{j}\|},\ \frac{(x_{i}-\overline{x})\times(x_{j}-\overline{x})}{\|(x_{i}-\overline{x})\times(x_{j}-\overline{x})\|},\ \frac{(x_{i}-x_{j})\times\left((x_{i}-\overline{x})\times(x_{j}-\overline{x})\right)}{\|(x_{i}-x_{j})\times\left((x_{i}-\overline{x})\times(x_{j}-\overline{x})\right)\|}\right)$
@@
 Efficient and expressive equivariant message-passing neural networks.
-Following (Du et al., 2024), we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames.
+Following [18], we leverage an efficient and expressive extension of the above equivariant message-passing neural networks based on local frames.
 We consider two modules to improve the expressiveness while maintaining the efficiency of the network.
@@
 Rotary positional embedding as invariant frame transition.
-Positional embedding are known to improve and stabilize the training of transformers (Shaw et al., 2018).
-We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture (Su et al., 2024).
+Positional embedding are known to improve and stabilize the training of transformers [48].
+We introduce the elegant relationship between the commonly used Rotary Position Embedding (RoPE) and frame transition we use in our architecture [52].
 In fact the frame transition can be considered as a generalized rotary positional embedding for 3D equivariant frames.
 Note that the spirit of rotary position embedding is to design an embedding scheme for relative position in sequential data.
-To build the rotary positional embedding,  (Su et al., 2024) utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
+To build the rotary positional embedding,  [52] utilizes the relative index between nodes $x_{i}$ and $x_{j}$ of a one-dimensional sequence to find a function $g(x_{i},x_{j},i-j)$ such that
 $\langle f_{q}(x_{i},i),f_{k}(x_{j},j)\rangle=g(x_{i},x_{j},i-j)$
 where $f_{q}$ and $f_{k}$ denote the neural network embeddings of query and key inside an attention layer.
-Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices (Su et al., 2024):
+Fortunately, there is an explicit way of building $g$ by multiplying $f_{\{q,k\}}(x_{i},i)$ with the diagonal matrix expanded by rotation matrices [52]:
 $\begin{pmatrix}\cos\theta_{n},&-\sin\theta_{n}\\ \sin\theta_{n},&\cos\theta_{n}\end{pmatrix}$
@@
 Temporal connection.
-In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network (Chen et al., 2018).
+In addition to commonly adapted message passing in the spatial domain, we consider another dimension, temporal domain, which can be described by the depth of the neural network [12].
 As message-passing neural networks are local, we believe it is beneficial to integrate multi-scale information.
@@
 […]
-Supplementary Information
-AlphaNet: Scaling Up Local Frame-based NNIP
+Supplementary Information for
+AlphaNet: Scaling Up Local-frame-based Neural Network Interatomic Potentials
 Appendix A Local Structures, Completeness, and Expressive Power of GNNs
@@
 These substructures define distinct local isomorphism types: (1) Tree isomorphism, where subgraphs share identical hierarchical parent-child relationships; (2) Triangular isomorphism, requiring matching cyclic triplets (e.g., molecular bond angles); and (3) Subgraph isomorphism, preserving node features and connectivity patterns.
-The expressive power of permutation-invariant GNNs is classically evaluated via the 1-Weisfeiler-Lehman (1-WL) test  (Xu et al., 2019).
-As established by  (Wijesinghe and Wang, 2022), the 1-WL test is equivalent to a GNN’s capacity to discriminate local subtree isomorphism, measuring its ability to distinguish hierarchical neighborhood structures.
+The expressive power of permutation-invariant GNNs is classically evaluated via the 1-Weisfeiler-Lehman (1-WL) test  [59].
+As established by  [58], the 1-WL test is equivalent to a GNN’s capacity to discriminate local subtree isomorphism, measuring its ability to distinguish hierarchical neighborhood structures.
 However, this framework operates solely on 2D graph topologies, ignoring 3D geometric relationships (e.g., bond angles, torsional dihedrals), thereby failing to address molecular or material systems where spatial configurations are decisive.
@@
 We can extends completeness to local geometric isomorphism, where isomorphism is defined under 3D geometric constraints (distances, angles, chiralities).
-Concurrently, methods encoding 3D geometric invariants  (Wijesinghe and Wang, 2022) have demonstrated enhanced expressivity, enabling differentiation of graphs with distinct local atomic environments beyond WL limitations.
-Similar to LeftNet (Du et al., 2023), AlphaNet is able to differentiate local structures beyond tree isomorphism by defining the 3D structure weights.
+Concurrently, methods encoding 3D geometric invariants  [58] have demonstrated enhanced expressivity, enabling differentiation of graphs with distinct local atomic environments beyond WL limitations.
+Similar to LeftNet [17], AlphaNet is able to differentiate local structures beyond tree isomorphism by defining the 3D structure weights.
 Appendix B Experimental Details
 B.1 Datasets
-The Defected Bilayer Graphene Dataset (Ying et al., 2024) consists of reference structures specifically designed to train and validate machine learning potentials (MLPs).
+The Defected Bilayer Graphene Dataset [61] consists of reference structures specifically designed to train and validate machine learning potentials (MLPs).
 It includes three bilayer systems: V0V0 (pristine), V0V1 (single vacancy on the top layer), and V1V1 (single vacancy in both layers).
@@
 Farthest point sampling (FPS) and principal component analysis (PCA) were used to ensure a representative and balanced division.
-The Formate Decomposition on Cu Dataset (Batzner et al., 2022) consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond.
+The Formate Decomposition on Cu Dataset [5] consists of configurations related to the decomposition process of formate on a Cu surface, specifically involving the cleavage of the C-H bond.
 The dataset includes initial, intermediate, and final states, such as monodentate and bidentate formate on Cu <110>, as well as the final state with an H ad-atom and desorbed CO2 in the gas phase.
-Nudged elastic band (NEB) method (Henkelman et al., 2000) was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations.
+Nudged elastic band (NEB) method [26] was used to generate an initial reaction path for the C-H bond breaking, followed by 12 short ab initio molecular dynamics (AIMD) simulations.
 These simulations were performed using the CP2K code, resulting in a total of 6855 DFT structures with a time step of 0.5 fs and 500 steps per trajectory.
@@
 Zeolites are highly valued porous materials that have been the subject of extensive scientific research due to their broad applicability.
-The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) (Car and Parrinello, 1985) simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) (Kresse and Furthmüller, 1996; Kresse and Hafner, 1993).
+The dataset includes 16 distinct zeolites, with atomic-level trajectories obtained through Ab Initio Molecular Dynamics (AIMD) [10] simulations conducted at 2000 K using the Vienna Ab Initio Simulation Package (VASP) [33, 34].
 For each zeolite, 80,000 snapshots were extracted, and the corresponding energies and atomic forces were calculated for every configuration.
 The dataset was randomly divided into training, validation, and test sets in a 6:2:2 ratio, comprising 48,000, 16,000, and 16,000 structures, respectively.
-The OC2M Dataset (Chanussot et al., 2021) is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials.
+The OC2M Dataset [11] is a large-scale subset of the OC20 Dataset, specifically designed for training and evaluating machine learning models in the context of interatomic potentials.
 This dataset includes approximately 2 million atomic configurations in the training set, with validation and test sets containing 4 splits, which is In Domain (ID), out-of-domain adsorbate (OOD adsorbate), out-of-domain catalyst (OOD catalyst), and both of unseen adsorbate and unseen catalyst (OOD both).
@@
 This extensive dataset is particularly valuable for developing and testing machine learning models that require a diverse and large-scale collection of training examples to achieve robust generalization across different types of atomic interactions.
-The MPTraj Dataset and Matbench Discovery Test (Riebesell et al., 2024).
-MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials (Deng et al., 2023).
-Matbench Discovery is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set (Wang et al., 2021).
+The MPTraj Dataset and Matbench Discovery Test [44].
+MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials [15].
+Matbench Discovery is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set [57].
 In this case we train our model on the MPTraj dataset and test it on the WBM test set.
@@
 Learning rate (lr) 5e-4 5e-4 5e-4 1e-4 1e-4
+grad_clip 0.5 0.5 0.5 0.5 0.5
 Scheduler Cos. Cos. Cos. Cos. Cos.
@@
 For computational modeling configurations, NequIP achieves optimal performance with tensor rank 2, feature size 32, and 5 Åcutoff in the formate decomposition dataset, while using tensor rank 1 with 7 Åcutoff for defective graphene studies.
-The zeolite dataset employs a fine-tuned DPA-2 potential via Deep Pot implementation ((Zhang et al., 2024b)).
-Models trained on OC20 leverage pre-trained SchNet/DimeNet++ architectures from https://fair-chem.github.io/core/model_checkpoints.html, with other results cross-validated against (Qu and Krishnapriyan, 2024) using publicly accessible validation sets.
-MPtrj benchmark data integrates results from https://matbench-discovery.materialsproject.org/ and comparative analysis in (Qu and Krishnapriyan, 2024).
+The zeolite dataset employs a fine-tuned DPA-2 potential via Deep Pot implementation ([63]).
+Models trained on OC20 leverage pre-trained SchNet/DimeNet++ architectures from https://fair-chem.github.io/core/model_checkpoints.html, with other results cross-validated against [42] using publicly accessible validation sets.
+MPtrj benchmark data integrates results from https://matbench-discovery.materialsproject.org/ and comparative analysis in [42].
 Speed testing on zeolite structures maintains uniform 2-layer architectures with 5 Åcutoff, where PaiNN/SchNet employ 120-atom basis and 16 RBF basis, MACE uses tensor order 0-1 (feature size 16), NequIP extends to tensor order 0-2 (feature size 16), AlphaNet configures 32 hidden channels with 12 attention heads, and EquiformerV2 implements sphere/edge channels 16 with 16-dimensional attention hidden states and 2-head mechanisms.
@@
 Appendix D Data Scaling and Dynamic loss weights
-The experimental datasets were generated using VASP (Kresse and Furthmüller, 1996; Kresse and Hafner, 1993), where the energy per atom typically converges around $5\,\mathrm{eV}$ under our default training settings.
+The experimental datasets were generated using VASP [33, 34], where the energy per atom typically converges around $5\,\mathrm{eV}$ under our default training settings.
 These configurations demonstrated robust performance within this energy regime.
-However, when applied to systems with significantly higher energy magnitudes per atom and greater energy variance—such as those produced by CP2K (Kühne et al., 2020) simulations—the original settings failed to achieve stable convergence.
+However, when applied to systems with significantly higher energy magnitudes per atom and greater energy variance—such as those produced by CP2K [35] simulations—the original settings failed to achieve stable convergence.
 To address this challenge, we developed a dynamic optimization strategy: initially adjusting the energy-to-force loss weight ratio from $\lambda_{E}:\lambda_{F}=4:100$ to $0.05:100$, then progressively increasing $\lambda_{E}$ while simultaneously decaying the learning rate.